### v1

# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated strong reasoning and memorization capabilities via pretraining on massive textual corpora. However, training LLMs on human-written text entails significant risk of privacy and copyright violations, which demands an efficient machine unlearning framework to remove knowledge of sensitive data without retraining the model from scratch. While Gradient Ascent (GA) is widely used for unlearning by reducing the likelihood of generating unwanted information, the unboundedness of increasing the cross-entropy loss causes not only unstable optimization, but also catastrophic forgetting of knowledge that needs to be retained. We also discover its joint application under low-rank adaptation results in significantly suboptimal computational cost vs. generative performance trade-offs. In light of this limitation, we propose two novel techniques for robust and cost-efficient unlearning on LLMs. We first design an Inverted Hinge loss that suppresses unwanted tokens by increasing the probability of the next most likely token, thereby retaining fluency and structure in language generation. We also propose to initialize low-rank adapter weights based on Fisher-weighted low-rank approximation, which induces faster unlearning and better knowledge retention by allowing model updates to be focused on parameters that are important in generating textual data we wish to remove. Through extensive experiments on the Training Data Extraction Challenge dataset using the GPT-Neo family of LLMs, we find that our approach can efficiently remove unwanted information with negligible deterioration in reasoning as well as generative capabilities.
1 Introduction
Recently, Large Language Models (LLMs) have demonstrated substantial performance gains in downstream tasks as their model size and the amount of training data increase (Zhao et al., 2023). This has prompted many research groups to prioritize the collection of more data for pretraining them and the development of more larger models than before (Brown et al., 2020; Chowdhery et al., 2023; Rae et al., 2021; Smith et al., 2022). However, this approach has introduced significant privacy concerns, as these models tend to memorize data indiscriminately Carlini et al. (2021; 2022). For example, Personally Identifiable Information (e.g., names, phone numbers, and email addresses) can be easily extracted from the LLMs (Carlini et al., 2021). Additionally, OpenAI is currently facing several lawsuits related to alleged copyright infringement within their pretraining text corpus (Grynbaum & Mac, 2023). In response to these challenges and in light of the recent GDPR legislation on the right to be forgotten (Voigt & Von dem Bussche, 2017; Rosen, 2011; Villaronga et al., 2018), machine unlearning for LLMs has emerged as a critical and rapidly growing research field (Yao et al., 2023; Si et al., 2023a).
Exact unlearning can be achieved by filtering out sensitive information and retraining from scratch. However, this process is highly resource-intensive and inefficient, especially when data deletion requests occur serially rather than in large batches. As shown in Figure 1, this necessitates the development of approximate unlearning frameworks, where the goal is to remove knowledge of specific data instances from a pretrained model. Consequently, several methods for knowledge unlearning have been proposed (Wang et al., 2023; Chen & Yang, 2023; Jang et al., 2023). Among them, Jang et al. (2023) proposed a simple approach that fine-tunes LLMs using Gradient Ascent (GA) on the forgetting data, while also introducing several metrics to assess its effectiveness. They demonstrate that GA not only can be used for successful knowledge unlearning but also minimizes catastrophic forgetting of other knowledge within the LLMs. Nevertheless, GA still has significant drawbacks, such as the issue of training not converging due to unbounded loss and high unlearning cost caused by full fine-tuning for the entire LLM model.
Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it lose knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion. After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other data points in the retain set ${\mathcal{D}}_{r}$ as well as preserve its previously acquired reasoning and generative capabilities.
On the other hand, LoRA (Hu et al., 2021) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning for downstream tasks. The core idea behind LoRA is to freeze the pretrained LLM’s weights and train rank decomposition matrices within each layer during adaptation to downstream tasks. This approach not only significantly reduces the number of trainable parameters but also achieves performance comparable to full fine-tuning. Recently, Biderman et al. (2024) experimentally demonstrates that LoRA acts as a powerful regularization mechanism that prevents it from reaching the same performance levels as full fine-tuning. As a result, it results in less forgetting in retained knowledge. From this perspective, many of LoRA’s benefits are particularly advantageous in practical unlearning scenarios: 1) it significantly reduces the tuning cost for unlearning, and 2) it effectively prevents forgetting, making it a valuable method for achieving the important goal of retaining knowledge. Nonetheless, its application to LLM unlearning remains unexplored until now.
In this paper, we propose a novel approach for robust and efficient unlearning for LLMs by introducing a novel loss function and leveraging LoRA with a new initialization. First, we analyze the derivatives of the existing GA and highlight its shortcomings: 1) dispersed gradients causing unnecessary forgetting, and 2) an unbounded loss that fails to converge. To address these issues, we propose the Inverted Hinge loss (IHL). From the derivative analysis, we show that the IHL resolves all the problems of GA, enabling faster and robust unlearning. Second, we explore the intersection between LLM unlearning and cost-efficient fine-tuning using LoRA. We find that LoRA suffers from significantly suboptimal trade-offs between its downstream performance after unlearning and its ability to generate information requested for deletion. To achieve robust and efficient unlearning with LoRA, we introduce a new LoRA Initialization via Weighted Low-Rank Approximation. Extensive experiments on the LM-Eval and Dialogue benchmarks using Neo-GPT, Phi, and Llama2 LLMs show that our proposed Inverted Hinge Loss combined with LoRA achieves robust and cost-efficient unlearning than GA (See Figure 2).
Figure 2: Compute cost for successful unlearning vs. post-unlearning downstream performances of GPT-Neo-125M. Each point represents a different forget set and LoRA rank (if used). Left: Accuracy averaged across 9 classification tasks (higher is better). Middle: F1 score averaged across 4 dialogue generation tasks (higher is better). Right: Perplexity on the validation set of the Pile dataset (lower is better). Dashed lines indicate the performances of the model prior to unlearning. Unlearning with vanilla LoRA leads to significant loss in performance compared to full-parameter unlearning. However, our proposed method using both the Inverted Hinge Loss and FLoRA performs competitively to full-unlearning in all three aspects while enjoying high parameter- and cost-efficiency of LoRA.
2 Related Work
Machine Unlearning. The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pre-trained model to discard information acquired from a specific subset of data. Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2021; Mehta et al., 2022; Chundawat et al., 2022; Cha et al., 2024). Recently, its significance has grown notably in Large Language Models (LLMs) due to its pivotal role in managing unintended memorization intrinsic to LLMs (Si et al., 2023b; Yao et al., 2024).
First, Wang et al. (2023) introduces the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets. Chen & Yang (2023) proposes the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model frozen. While these approaches demonstrates impressive unlearning results across various datasets, they are hindered by significant drawbacks, such as the necessity to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) and the need to finetune models using them (Wang et al., 2023). Parameter merging-based methods recently garner attention as another category of unlearning for LLMs (Wu et al., 2023; Ilharco et al., 2022). Experimental results demonstrate that simple task arithmetic can facilitate cost-efficient unlearning for LLMs. However, these methods also exhibit limitations, such as achieving a relatively weak level of unlearning (Ilharco et al., 2022) or incurring higher computational costs to detect privacy neurons (Wu et al., 2023).
In contrast, following an unlearning approach similar to that used in image classification (Golatkar et al., 2020; Cha et al., 2024), Jang et al. (2023) leverages Gradient Ascent (GA) to maximize the next-token prediction loss for a sequence of tokens from the forget data. Unlike in image classification, they demonstrate that LLMs do not suffer from forgetting of retained knowledge when using GA on the forget data. Consequently, they show that effective unlearning for LLMs is achievable solely using the forget dataset, while preserving general LLM knowledge and maintaining performance on downstream tasks.
Parameter-Efficient Fine-Tuning. Fine-tuning pretrained LLMs to adapt towards specific downstream tasks and instructions comes at an intractable computational cost due to the large model size and complexity. To alleviate this burden, previous work has developed Parameter-Efficient Fine-Tuning (PEFT) methods that adapt only a small portion of LLM parameters while freezing the other pretrained parameters intact (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023). Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2021; Zhang et al., 2023; YEH et al., 2023; Kopiczko et al., 2023; Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter. A key benefit of LoRA is that the adapter weights can be merged seamlessly to pretrained parameters after fine-tuning, such that the post-adaptation LLM shares the same inference cost as the base pretrained LLM. While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposes to initialize LoRA weights using the principal singular vectors and values of the linear weights. Inspired by this work, we propose to extract Fisher information from the data requested for unlearning and perform weighted low-rank approximation to prepare a LoRA initialization that induces faster unlearning and better preservation of general knowledge
3 Proposed Method
In this section, we introduce the notations and provide preliminaries on unlearning through Gradient Ascent (GA) and low-rank adaptation (LoRA). We then analyze the gradient properties of GA and highlight its inherent issues. Building on this analysis, we propose the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA. Finally, we present the idea of employing LoRA with a novel initialization strategy, thereby enabling cost-efficient unlearning.
3.1 Preliminaries
Problem and notation. Given a sequence of $T$ tokens ${\bm{x}}=(x_{1},x_{2},\dots,x_{T})$, a language model (LM) models the likelihood of the sequence via next token prediction: $p_{\theta}({\bm{x}})=\prod_{t=1}^{T}p_{\theta}(x_{t}|x_{<t})$. After pretraining, we assume that an end-user has requested to delete a subset of the training set $\mathcal{D}_{f}\subset\mathcal{D}$, which we refer to as the forget set. Also, we consider a retain set $\mathcal{D}_{r}$ that represents the auxilary dataset which contains general knowledge the model should not forget after unlearning (e.g., Wikitext dataset (Merity et al., 2016)).
Gradient Ascent. Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, which leads to a simple yet effective baseline of gradient ascent (GA) (Cha et al., 2024; Jang et al., 2023). Unlike the usual gradient descent, GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
$\mathcal{L}_{GA}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$
Note that in practice, the log-likelihood is implemented in terms of cross-entropy losses on incremental tokens. Thus, GA tunes the model towards maximizing the cross-entropy, the value of which is unbounded above by nature, making optimization ill-posed and unstable.
Low-Rank Adaptation. Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank, LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter. In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes
$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}$ (1)
During fine-tuning, the original weight ${\bm{W}}$ is kept frozen and only the low-rank factors ${\bm{A}}$ and ${\bm{B}}$ are updated via gradient descent. To ensure that the initial attachment of LoRA adapters does not alter the output of the LLM, LoRA defaults to initializing ${\bm{A}}$ with a Kaiming-uniform distribution (He et al., 2015) and ${\bm{B}}$ as the zero matrix. Note that after fine-tuning, LoRA adapters can simply be merged with the original weights ${\bm{W}}^{\prime}={\bm{W}}+{\bm{B}}{\bm{A}}$, thereby avoiding any additional latency during inference. In the context of unlearning, this feature of LoRA naturally promotes a layer of security as it prevents adversarial agents from tracking previous records of unlearning just by inspecting the model architecture.
3.2 Preliminary Results
Despite its wide use in domain adaptation and instruction tuning, LoRA is not yet explored under the task of LLM unlearning to the best of our knowledge. Therefore, we first share empirical results from low-rank adapting LLMs using gradient difference (Maini et al., 2024; Liu et al., 2022a) as our objective to motivate our approach. For this experiment, we follow previous work (Jang et al., 2023) unlearn 32 text sequences from the Training Data Extraction Challenge from GPT-Neo-1.3B model pretrained on the Pile dataset (Gao et al., 2020). More information on the details on the setup can be found in Section 4 and the Appendix.
Figure 2 shows the results. Notably, vanilla LoRA suffers from lack of plasticity and ends up failing to sufficiently unlearn ${\mathcal{D}}_{f}$ within 20 epochs. When running more unlearning epochs or increasing the learning rate for sufficient unlearning, the model loses its previously acquired reasoning and generative capabilities, as shown in the significant decrease in LM-eval and Dialogue performances. In the remainder of this section, we present two techniques towards making LLM unlearning viable while enjoying the compute-efficiency of LoRA.
3.3 Inverted Hinge Loss: Novel Loss Function for Unlearning in LLMs
Motivation: Inherent Limitation of GA. We analyze the inherent issues of GA from the perspective of its gradients. The output layer of a language model is a softmax layer that outputs probabilities over the vocabulary. Let ${y}_{t}$ be the logits (pre-softmax activations) produced by the LLM model for the $t$-th token, and let $V$ be the vocabulary size. The probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function: $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{\sum_{v=1}^{V}\exp(y_{t}^{(v)})}$ where $y_{t}^{(x_{t})}$is the logit corresponding to the true token $x_{t}$ and $y_{t}^{(v)}$ is the logit corresponding to the $v$-th token in the vocabulary. When we use $\mathcal{L}_{GA}$ for unlearning for LLMs, the gradient of the log-probability with respect to the logits is:
$\frac{\partial\log(p_{\theta}(x_{t}|x_{<t}))}{\partial y_{t}^{(v)}}=\begin{cases}1-p_{\theta}(x_{t}|x_{<t})&\text{if }v=x_{t}\\ -p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
Given this, the derivative of $\mathcal{L}_{GA}({\bm{x}})$ with respect to the logits at each $t$ becomes $\nabla_{y_{t}^{(v)}}\mathcal{L}_{GA}({\bm{x}})={\partial\log(p_{\theta}(x_{t}|x_{<t}))}/{\partial y_{t}^{(v)}}$. From this derivative of GA, we can confirm its unlearning mechanism and inherent issues. First, when $v=x_{t}$, GA decreases the prediction score corresponding to the true token while increasing the prediction score for all other tokens. This demonstrates that GA performs unlearning by encouraging the model to predict tokens different from the true token.
However, despite the successful experimental results of GA (Jang et al., 2023) in LLMs, our gradient analysis reveals its inherent issues. When $v=x_{t}$, it increases the prediction scores of all tokens except the true token. As a result, GA has problems: 1) the loss not converging and becoming unbounded, and 2) unnecessary additional forgetting due to the increase in logits for all other tokens, 3) a large vocabulary size induces gradients to spread out across all other logits, making gradient updates for unlearning inefficient.
Inverted Hinge Loss (IHL). Considering the limitations of GA discussed above, our motivation is to develop a simple yet novel loss function for unlearning in LLMs. The new loss function aims to effectively achieve unlearning by decreasing the prediction score of the true token while increasing the prediction scores of only a minimal number of other tokens. Inspired by the hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge Loss (IHL) as shown below:
$\mathcal{L}_{\text{IHL}}({\bm{x}})=\max\left(0,1+p_{\theta}(x_{t}|x_{<t})-\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))\right)$
Considering the probability $p_{\theta}(x_{t}|x_{<t})$ given by the softmax function, the derivative of $\mathcal{L}_{\text{IH}}({\bm{x}})$ with respect to $y_{t}^{(v)}$ is:
$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}}.\end{cases}$
Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IH}}({\bm{x}})\neq 0$. Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IH}}({\bm{x}})=0$. The detailed process for deriving this is provided in the Appendix A
The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly demonstrates how the Inverted Hinge loss addresses the shortcomings of GA in knowledge unlearning for LLMs. First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with the opposite sign). This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly and intensively but also prevents spreading out of gradients for unlearning. During this process, the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ increase slowly in proportion to the difference between the predictions for $x_{t}$ and $v^{\star}$. Second, once knowledge unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$), the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease. This not only prevents unnecessary forgetting but also results in a bounded loss form.
3.4 FLoRA: Novel LoRA Initialization for LLM Unlearning
Motivation. Aside from reduing the number of trainable parameters and thus reducing the memory requirements of LLM fine-tuning, LoRA serves as a potent form of regularization. While it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios (Biderman et al., 2024). Under the task of LLM unlearning, however, our objective strictly requires to completely remove knowledge of ${\mathcal{D}}_{f}$, and the strong stability from LoRA imposes a strong burden: vanilla LoRA unlearning requires large number of iterations through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in our Figure 2). Inspired by previous work on PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are important to ${\mathcal{D}}_{f}$ but not so important to ${\mathcal{D}}_{r}$ would benefit the unlearning process in terms of both convergence and post-unlearning performance. We hence design a novel LoRA initialization technique that measures the Fisher information on each weight ${\bm{W}}$ targeted for LoRA, and perform weighted low-rank decomposition from which we use the principal components to initialize adapter weights ${\bm{A}}$ and ${\bm{B}}$.
Parameter Importances via Fisher Information. The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information dataset ${\mathcal{D}}$ provides on model parameters $\theta$. More concretely, ${\bm{F}}_{\theta}$ is computed as the second cross-moments of first partial derivatives of the log-likelihood of ${\mathcal{D}}$ (left of Eq. 3.4). However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead which, in the context of LLMs, can be computed as the average of squared gradients propagated from the language modeling objective (i.e. cross entropy loss):
${\bm{F}}_{\theta}=\mathbb{E}_{\mathcal{D}}\left[\left(\dfrac{\partial}{\partial\theta}\log p{\color[rgb]{0,0,0}{}_{\theta}}({\mathcal{D}}|\theta)\right)^{2}\right]\approx\dfrac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}\in{\mathcal{D}}}\left(\dfrac{\partial}{\partial\theta}\mathcal{L}_{\text{LM}}({\bm{x}};\theta)\right)^{2}\eqcolon\hat{{\bm{F}}}_{\theta},$ (2)
where $\mathcal{L}_{\text{LM}}$ is the next-token prediction loss of LM, such as $\mathcal{L}_{\text{LM}}({\bm{x}};\theta)=\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$. Within our LLM unlearning setup, a high empirical Fisher information measured with ${\mathcal{D}}_{f}$ indicates that the next-token prediction loss on ${\mathcal{D}}_{f}$ leads to large absolute gradients on the parameter under concern, and we consider such parameters to be important in generating sequences in ${\mathcal{D}}_{f}$.
Let $\hat{{\bm{F}}}_{\bm{W}}^{f}$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp. $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$). Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{rel}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to first identify parameters that are important specifically for ${\mathcal{D}}_{f}$, but unimportant to ${\mathcal{D}}_{r}$. While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus choose to use the relative Fisher information matrix rather than $\hat{{\bm{F}}}^{f}$.
Fisher-weighted LoRA Initialization (FLoRA). Intuitively, initializing LoRA weights to be parameters that are important in generating ${\mathcal{D}}_{f}$ would benefit unlearning as gradients would only modify parameters important to ${\mathcal{D}}_{f}$ while keeping all other parameters intact. Given the relative importance $\hat{{\bm{F}}}_{\bm{W}}^{rel}$ for each target parameter ${\bm{W}}$, we propose to initialize LoRA adapter weights as the solution to the following problem of Weighted Low-Rank Approximation (WLRA):
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\sum_{i,j}\left([\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}({\bm{W}}-{\bm{B}}{\bm{A}})_{i,j}\right)^{2}$ (3)
Note that when all weights equal one, this problem reduces down to a standard low-rank approximation problem, of which the solution can easily be computed via rank-$r$ SVD. In general, however, this minimization problem does not have a closed-form solution and thus requires iterative optimization (Srebro & Jaakkola, 2003). While we may resort to iterative methods towards initializing LoRA weights in a fine-grained manner, this would undermine the gains in efficiency due to employing LoRA. Thus, we tackle a simplified version of WLRA instead by assuming that parameters in each row of ${\bm{W}}$ share the same importance, for which we use the square-root of the row-wise sum of the relative Fisher information matrix as
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\left\|\text{diag}\left((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\ \bm{\cdot}\ \mathds{1})^{\frac{1}{2}}\right)({\bm{W}}-{\bm{B}}{\bm{A}})\right\|_{2}$ (4)
where $\mathds{1}\in\mathbb{R}^{k}$ is the all-one vector. $\text{diag}(\cdot)$ and $\bm{\cdot}$ denote the vector diagonalization function and the matrix-vector product, respectively. Unlike general WLRA, the row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
Given the solution, we use the solution ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights. To ensure that the output is not perturbed after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}{\bm{A}}$. In summary, our Fisher-Weighted initialization extracts parameters that are important only to ${\mathcal{D}}_{f}$ and not to ${\mathcal{D}}_{r}$ such that LoRA tuning can be focused towards erasing knowledge of ${\mathcal{D}}_{f}$.
3.5 Final Loss Function for Unlearning
Following initialization with FLoRA, unlearning is performed on the model $\Theta=\theta\cup\theta_{FLoRA}$, which consists of the original weights $\theta$ and the LoRA weights for each layer $\theta_{FLoRA}=\{{\bm{A}}^{*}_{l},{\bm{B}}^{*}_{l}\}_{l=1}^{L}$, where $L$ represents the number of layers where FLoRA is applied. During unlearning process, $\theta$ remains frozen, and only updates to the FLoRA weights $\theta_{FLoRA}$ are made. Furthermore, when performing unlearning on the forget dataset $D_{f}$, we incorporate the auxiliary retain set $D_{r}$, which contains general knowledge (e.g., the Wikitext dataset), as a default setting, following the gradient difference approach (Liu et al., 2022a; Maini et al., 2024). The final loss function using the proposed Inverted Hinge Loss is defined as follows:
$\operatorname*{minimize}_{\theta_{FLoRA}}\sum_{{\bm{x}}_{r}\in D_{f},{\bm{x}}_{f}\in D_{r}}\mathcal{L}_{\text{IH}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (5)
In practice, training (unlearning) for the LLM model involves sampling corpora from each dataset, $D_{f}$ and $D_{r}$, and minimizing the final loss function through stochastic gradient descent.
4 Experiments
In this section, we first perform unlearning experiments based on the Training Data Extraction Challenge (§4.1), followed by ablation and analytical results (§4.2). All experiments were conducted on a remote GCP server equipped with 16 NVIDIA A100 Tensor Core GPUs.
Table 1: Evaluation results on LM-Evaluation, Dialogue datasets before and after unlearning samples from the TDEC dataset. Note that, in all experiments using LoRA, GD is used as the default loss function. The "+IHL" results refer to experiments where GA within GD is replaced with the proposed IHL (i.e., using Eqn.(5)), and "+FLoRA" indicates the results where FLoRA is used instead of LoRA.
Model Method Params. Epochs EL10 MA Reasoning Dialogue Pile
(%) (%$\downarrow$ (%$\downarrow$ (Acc)$\uparrow$ (F1$\uparrow$ (PPL)$\downarrow$
Before - - 30.9 77.4 43.4 9.4 17.8
GA 100.0 17.2 1.0 27.4 39.9 2.6 577.8
GPT-Neo GD 4.6 0.7 24.9 42.4 5.9 54.2
125M LoRA 1.6 8.6 0.3 20.6 40.8 2.5 129.4
+ IHL 11.4 0.4 22.7 41.9 6.0 32.9
+ FLoRA 6.0 0.3 23.9 42.2 10.1 24.0
Before - - 67.6 92.2 49.8 11.5 11.5
GA 100.0 13.8 1.9 30.4 49.7 8.5 15.8
GPT-Neo GD 12.8 2.2 30.9 48.4 12.7 10.8
1.3B LoRA 0.8 19.3 1.7 31.4 45.0 9.7 31.8
+ IHL 20.0 1.7 44.6 47.1 10.2 14.9
+ FLoRA 13.0 0.5 29.6 48.3 12.1 14.7
Before - - 70.4 93.4 52.3 11.5 10.4
GA 100.0 10.8 1.6 31.0 51.9 11.1 17.9
GPT-Neo GD 8.0 0.7 28.3 44.0 12.7 17.9
2.7B LoRA 0.7 14.0 0.1 20.4 45.9 6.7 61.1
+ IHL 17.8 0.0 26.7 49.6 8.5 22.2
+ FLoRA 10.3 0.1 28.5 49.6 10.7 16.0
4.1 Training Data Extraction Challenge
Experimental Setup.
Assuming a scenario where an adversary attacks a pretrained LLM to extract text samples previously used in training, the Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) that are found to be easily extractable. For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$. For the retain set ${\mathcal{D}}_{r}$, we use the validation subset of WikiText as it contains factual world knowledge that we wish to maintain within the LLM.
For TDEC experiments, we consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using three different forget sets. For this experiment, we use a fixed learning rate of 5e-5 when full finetuning, and 2e-4 when using LoRA and FLoRA. For all model sizes, we use a LoRA (and FLoRA) rank of $r=16$ as a default. We set the default to applying LoRA to the attention matrices for Query (Q) and Value (V), as well as to the Feedforward Network (FFN).
Following previous work (Jang et al., 2023), we measure the unlearning efficacy via two metrics. The ${\bm{n}}$-gram Extraction Likelihood (ELn) measures the $n$-gram overlap between the ground truth sequence in ${\mathcal{D}}_{f}$ and the output generated by the model. The Memorization Accuracy (MA) measures the token-wise accuracy of the LLM on ${\mathcal{D}}_{f}$. More details on these metrics are introduced in Section B of the Appendix. After each unlearning epoch, we measure EL10/AM and consider the model to have successfully unlearned ${\mathcal{D}}_{f}$ if both values are smaller than those measured using a held-out validation set. Once unlearning is finished, we evaluate the unlearned model on various downstream benchmarks to ensure the LLM maintains its previously acquired language modeling capabilities after unlearning. To assess its reasoning capabilities, we average accuracies across 9 different classification datasets LAMBADA (Paperno et al., 2016), Hellaswag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), COPA (Gordon et al., 2012), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), PiQA (Bisk et al., 2020), MathQA (Amini et al., 2019), PubmedQA (Jin et al., 2019). To measure generative performance, we also measure the F1 score averaged over four dialogue generation datasets, Blended Skill Talk (Smith et al., 2020), Empathetic Dialogues Rashkin et al. (2019), Wizard of Internet (Komeili et al., 2021), Wizard of Wikipedia (Dinan et al., 2018). Lastly, we measure the perplexity on the validation subset of the Pile (Gao et al., 2020).
Baseline.
For experiments on knowledge unlearning in LLMs, we select using two baseline algorithms, Gradient Ascent (GA) (Jang et al., 2023) and Gradient Difference (GD) (Liu et al., 2022a; Maini et al., 2024), which rely solely on the original language model while utilizing a forget dataset $D_{f}$ only or both a forget dataset $D_{f}$ and an auxiliary dataset $D_{r}$, respectively. For all experiments, we use the same forgetting criteria for ELn and MA as in  Jang et al. (2023) and report the results after performing unlearning until these criteria are met.
Figure 3: Results from unlearning 32 examples from Training Data Extraction Challenge. Each row represents the average performance in each set of LM capabilies: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better). Each column shows results under different model size: 125M, 1.3B, and 2.7B. Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed. X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics. Solid lines indicate the performance of different methods averaged only across successful unlearning trials. The dashed lines show respective performances prior to unlearning. Unlearning with vanilla GD not only leads to significant loss in performance, but also fails to unlearn sufficiently even with large LoRA ranks. Replacing the negative cross-entropy loss with our hinge loss (IHL) shows significant gains in reasoning and generation capabilities, but still fails to unlearn a number of example sets. Using NCE but with Fisher-weighted LoRA initialization (FLoRA) notably increases the rate of successful unlearning, but at the cost of significant loss in overall LM capability. Using both IHL and FLoRA best minimizes post-unlearning performance degradation in all three capabilities under successful unlearning.
Results.
Table 1 shows the evaluation results of full-parameter unlearning and unlearning with LoRa and FLoRA. From our experimental results, we can identify the following findings. First, for all model sizes, using GD for unlearning not only achieves the forgetting criteria in fewer epochs compared to using GA alone, but it also better preserves previously acquired knowledge. Specifically, for the 125M and 1.3B models, knowledge unlearning with GA leads to a significant deterioration in PPL on the Pile dataset, whereas GD not only overcomes this issue but also achieves improved Dialogue (F1) performance compared to Before (e.g., GPT-Neo 1.3B). Second, we observe that simply applying GD loss with LoRA fails to achieve successful unlearning across all model sizes. Despite the fact that only about 0.7% to 1.6% of the total parameters are updated for unlearning when using LoRA, its simple application actually causes forgetting of previously acquired knowledge. Finally, when using the proposed IHL along with FLoRA, it achieves the forgetting criteria in fewer epochs while better preserving previously acquired knowledge. Although using IHL with LoRA alone maintains previously acquired knowledge better than the baseline, it requires relatively more epochs to meet the forgetting criteria. However, when combined with FLoRA, not only is the number of epochs required to meet the forgetting criteria significantly reduced, but overall, previously acquired knowledge is preserved more successfully.
Additionally, we conduct further experiments with different ranks for LoRA and FLoRA, with results and detailed explanations provided in Figure 3 and its caption. From all experiments so far, we could confirm that simply applying LoRA to GD does not lead to successful unlearning. However, using the proposed IHL in combination with FLoRA demonstrates better cost-efficiency than the existing GA and GD methods, while also better preserving previously acquired knowledge during unlearning.
Figure 4: Results from unlearning 32 examples from Training Data Extraction Challenge using LoRA for GPT-Neo 2.7B. Note that each color denotes Unlearning with vanilla GD, Replacing the negative cross-entropy loss with our hinge loss (IHL), Using NCE but with Fisher-weighted LoRA initialization (FLoRA), and Using both IHL and FLoRA. Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed. X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics.
4.2 Analysis
What modules do we need to adapt?
Figure 4 presents experiments where LoRA or FLoRA is applied to various weight matrices, including those for Query (Q), Value (V), Key (K), Output (O) in the self-attention module, and the FeedForward Network (FFN). While the original LoRA paper indicates that applying LoRA to Q and V yields superior performance on downstream tasks (Hu et al., 2021), our experiments show that applying LoRA to Q and V alone is insufficient to meet the unlearning criteria in all cases. Notably, applying LoRA to the FFN results in much more successful unlearning in cases of using FLoRA. Furthermore, applying FLoRA to QVFFN in conjunction with IHL not only achieves superior unlearning results but also better preserves previously acquired knowledge. Based on these findings, we adopt the application of LoRA and FLoRA to QVFFN as the default setting.
Cost-efficiency of the proposed method.
To compare unlearning costs, we evaluate the FLOPs for each method and the results are presented in Figure 2. While using vanilla LoRA offers significant advantages in terms of unlearning cost (i.e., FLOPs), as confirmed by our previous experiments, it results in substantial performance losses compared to full-parameter unlearning. In contrast, combining the proposed Inverted Hinge Loss with FLoRA not only achieves the best performance but also leverages the cost advantages of LoRA, demonstrating that this approach enables cost-efficient unlearning.
5 Concluding Remarks
In this paper, we address limitations of existing unlearning techniques (i.e., Gradient Ascent) by introducing the novel Inverted Hinge Loss (IHL) and leveraging LoRA with a novel initialization (i.e., Fisher-weighted LoRA Initialization (FLoRA)). Our findings reveal that IHL overcomes key drawbacks associated with Gradient Ascent (GA), such as dispersed gradients and unbounded loss, leading to faster and more robust unlearning. Furthermore, our approach of combining FLoRA with IHL not only reduces the cost of unlearning but also enhances the preservation of previously acquired knowledge. Through experiments, we show that our method achieves superior performance in knowledge unlearning compared to the baseline method.
Appendix A Derivative Analysis for the Inverted Hinge Loss Function
The function $p_{\theta}(x_{t}|x_{<t})$ represents a probability distribution that indicates the likelihood of $x_{t}$ taking a specific token $x_{t}$ given the previous tokes $x_{<t}$. This probability is expressed using the softmax function: $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{\sum_{v=1}^{V}\exp(y_{t}^{(v)})}$, where $y_{t}^{(v)}$ denotes the score for the $v$-th token in the vocabulary. To differentiate this function with respect to $y_{t}^{(x_{t})}$, we rewrite $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{Z}$ where $Z=\sum_{v=1}^{V}\exp(y_{t}^{(v)})$ is the normalization constant.
We differentiate this function with respect to $y_{t}^{(k)}$ considering two cases: 1) $k=x_{t}$ and 2) $k\neq x_{t}$. For the first case, we can get the following by using the chain rule:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(x_{t})}}=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(\frac{\exp(y_{t}^{(x_{t})})}{Z}\right)=\frac{1}{Z}\frac{\partial\exp(y_{t}^{(x_{t})})}{\partial y_{t}^{(x_{t})}}-\frac{\exp(y_{t}^{(x_{t})})}{Z^{2}}\frac{\partial Z}{\partial y_{t}^{(x_{t})}}$
Here, $\frac{\partial\exp(y_{t}^{(x_{t})})}{\partial y_{t}^{(x_{t})}}=\exp(y_{t}^{(x_{t})})$ and $\frac{\partial Z}{\partial y_{t}^{(x_{t})}}=\exp(y_{t}^{(x_{t})})$. Therefore, it becomes:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(x_{t})}}=\frac{\exp(y_{t}^{(x_{t})})}{Z}-\frac{\exp(y_{t}^{(x_{t})})^{2}}{Z^{2}}=p_{\theta}(x_{t}|x_{<t})-p_{\theta}(x_{t}|x_{<t})^{2}=p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))$
For the first case, using the chain rule again, we get:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(k)}}=\frac{\partial}{\partial y_{t}^{(k)}}\left(\frac{\exp(y_{t}^{(x_{t})})}{Z}\right)=-\frac{\exp(y_{t}^{(x_{t})})}{Z^{2}}\frac{\partial Z}{\partial y_{t}^{(k)}}$
where $\frac{\partial Z}{\partial y_{t}^{(k)}}=\exp(y_{t}^{(k)})$. Therefore,
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(k)}}=-\frac{\exp(y_{t}^{(x_{t})})\exp(y_{t}^{(k)})}{Z^{2}}=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(k|x_{<t})$
Thus, we can summarize them as below:
$\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))&\text{if }v=x_{t}\\ -p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
Based on the derivative of $p_{\theta}(x_{t}|x_{<t})$ above , we can calculate the derivative of $\mathcal{L}_{\text{IH}}$. Firstly, for convenience, we define $p_{t}=p_{\theta}(x_{t}|x_{<t})$ and $\hat{p}_{t}=\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))$. The loss function can be rewritten as:
$\mathcal{L}_{\text{IH}}({\bm{x}})=\max\left(0,1+p_{t}-\hat{p}_{t}\right)$
To calculate the derivative of $\mathcal{L}_{\text{IH}}$, we need to consider three cases: 1) when $v=x_{t}$, 2) when $v=v^{\star}$ where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$, 3) when $v\neq x_{t}$ and $v\neq v^{\star}$. In the case where $01+p_{t}-\hat{p}_{t}>$ 0, using the derivative of $p_{\theta}(x_{t}|x_{<t})$ mentioned earlier, the derivative of $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ is as follows:
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(x_{t})}}$ $=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
$=p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))+p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v^{\star}|x_{<t})$
$=p_{\theta}(x_{t}|x_{<t})\left(1-p_{\theta}(x_{t}|x_{<t})+p_{\theta}(v^{\star}|x_{<t})\right)$
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v^{\star})}}$ $=\frac{\partial}{\partial y_{t}^{(v^{\star})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
$=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})(1-p_{\theta}(v^{\star}|x_{<t}))$
$=-p_{\theta}(v^{\star}|x_{<t})\left(1-p_{\theta}(v^{\star}|x_{<t})+p_{\theta}(x_{t}|x_{<t})\right)$
$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v)}}$ $=\frac{\partial}{\partial y_{t}^{(v)}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
$=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v|x_{<t})+p_{\theta}(v^{\star}|x_{<t})\cdot p_{\theta}(v|x_{<t})$
$=p_{\theta}(v|x_{<t})\left(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})\right)$
Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $1+p_{t}-\hat{p}_{t}\leq 0$. In summary, the derivatives of the loss function $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ for the three cases are:
$\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$
Appendix B Evaluation Metrics
How to measure success of unlearning? Following previous work Jang et al. (2023); Tirumala et al. (2022), we empirically measure the success of unlearning using two metrics, Extraction Likelihood (EL) and Memorization Accuracy (MA), which we briefly discuss below.
After unlearning each sequence ${\bm{x}}=(x_{1},\dots,x_{T})\in\mathcal{D}_{f}$, the Extraction Likelihood (EL) is measured as the $n$-gram overlap between the ground truth sequence ${\bm{x}}$ and the output of the model after unlearning.
$\textsc{Overlap}_{n}({\bm{a}},{\bm{b}})=\dfrac{\sum_{{\bm{c}}\in\textsc{$n$-gram}({\bm{a}})}\mathds{1}\{{\bm{c}}\in\textsc{$n$-gram}({\bm{b}})\}}{|\textsc{$n$-gram}({\bm{a}})|}$ (6)
$\textsc{EL}_{n}({\bm{x}})=\dfrac{\sum_{t=1}^{T-n}\textsc{Overlap}_{n}\left(f_{\theta}(x_{<t}),x_{\geq t}\right)}{T-n}$ (7)
The Memorization Accuracy (MA) measures the token-wise memorization of the LM $p_{\theta}$.
$\textsc{MA}({\bm{x}})=\dfrac{\sum_{t=1}^{T}\mathds{1}\{\operatorname*{arg\,max}_{x}p_{\theta}(x|x_{<t})=x_{t}\}}{T-1}$ (8)
Given these two metrics, we flag successful unlearning when the average EL and MA on $\mathcal{D}_{f}$ goes below the EL and MA values measured on the validation set unseen during training. In our experiments we measure EL with 10-grams, which results in the following early stopping criterion.
$\dfrac{1}{\mathcal{D}_{f}}\sum_{{\bm{x}}\in\mathcal{D}_{f}}\textsc{EL}_{10}({\bm{x}})\leq\dfrac{1}{\mathcal{D}_{\text{val}}}\sum_{{\bm{x}}\in\mathcal{D}_{\text{val}}}\textsc{EL}_{10}({\bm{x}})\;\;\;\;\text{and}\;\;\;\;\dfrac{1}{\mathcal{D}_{f}}\sum_{{\bm{x}}\in\mathcal{D}_{f}}\textsc{MA}({\bm{x}})\leq\dfrac{1}{\mathcal{D}_{\text{val}}}\sum_{{\bm{x}}\in\mathcal{D}_{\text{val}}}\textsc{MA}({\bm{x}})$

### v2

@@
 Large Language Models (LLMs) have demonstrated strong reasoning and memorization capabilities via pretraining on massive textual corpora.
-However, training LLMs on human-written text entails significant risk of privacy and copyright violations, which demands an efficient machine unlearning framework to remove knowledge of sensitive data without retraining the model from scratch.
-While Gradient Ascent (GA) is widely used for unlearning by reducing the likelihood of generating unwanted information, the unboundedness of increasing the cross-entropy loss causes not only unstable optimization, but also catastrophic forgetting of knowledge that needs to be retained.
-We also discover its joint application under low-rank adaptation results in significantly suboptimal computational cost vs.
-generative performance trade-offs.
-In light of this limitation, we propose two novel techniques for robust and cost-efficient unlearning on LLMs.
-We first design an Inverted Hinge loss that suppresses unwanted tokens by increasing the probability of the next most likely token, thereby retaining fluency and structure in language generation.
-We also propose to initialize low-rank adapter weights based on Fisher-weighted low-rank approximation, which induces faster unlearning and better knowledge retention by allowing model updates to be focused on parameters that are important in generating textual data we wish to remove.
-Through extensive experiments on the Training Data Extraction Challenge dataset using the GPT-Neo family of LLMs, we find that our approach can efficiently remove unwanted information with negligible deterioration in reasoning as well as generative capabilities.
+However, this poses risk of privacy and copyright violations, highlighting the need for efficient machine unlearning methods that remove sensitive data without retraining from scratch.
+While Gradient Ascent (GA) is commonly used to unlearn by reducing the likelihood of generating unwanted content, it leads to unstable optimization and catastrophic forgetting of retrained knowledge.
+We also find that combining GA with low-rank adaptation results in poor trade-offs between computational cost and generative performance.
+To address these challenges, we propose two novel techniques for robust and efficient unlearning for LLMs.
+First, we introduce Inverted Hinge loss, which suppresses unwanted tokens while maintaining fluency by boosting the probability of the next most likely token.
+Second, we develop a data-adaptive initialization for LoRA adapters via low-rank approximation weighted with relative Fisher information, thereby focusing updates on parameters critical for removing targeted knowledge.
+Experiments on the Training Data Extraction Challenge dataset using GPT-Neo models as well as on the TOFU benchmark with Phi-1.5B and Llama2-7B models demonstrate that our approach effectively removes sensitive information while maintaining reasoning and generative capabilities with minimal impact.
 1 Introduction
-Recently, Large Language Models (LLMs) have demonstrated substantial performance gains in downstream tasks as their model size and the amount of training data increase (Zhao et al., 2023).
-This has prompted many research groups to prioritize the collection of more data for pretraining them and the development of more larger models than before (Brown et al., 2020; Chowdhery et al., 2023; Rae et al., 2021; Smith et al., 2022).
-However, this approach has introduced significant privacy concerns, as these models tend to memorize data indiscriminately Carlini et al.
-(2021; 2022).
-For example, Personally Identifiable Information (e.g., names, phone numbers, and email addresses) can be easily extracted from the LLMs (Carlini et al., 2021).
-Additionally, OpenAI is currently facing several lawsuits related to alleged copyright infringement within their pretraining text corpus (Grynbaum & Mac, 2023).
-In response to these challenges and in light of the recent GDPR legislation on the right to be forgotten (Voigt & Von dem Bussche, 2017; Rosen, 2011; Villaronga et al., 2018), machine unlearning for LLMs has emerged as a critical and rapidly growing research field (Yao et al., 2023; Si et al., 2023a).
-Exact unlearning can be achieved by filtering out sensitive information and retraining from scratch.
-However, this process is highly resource-intensive and inefficient, especially when data deletion requests occur serially rather than in large batches.
-As shown in Figure 1, this necessitates the development of approximate unlearning frameworks, where the goal is to remove knowledge of specific data instances from a pretrained model.
-Consequently, several methods for knowledge unlearning have been proposed (Wang et al., 2023; Chen & Yang, 2023; Jang et al., 2023).
-Among them, Jang et al.
-(2023) proposed a simple approach that fine-tunes LLMs using Gradient Ascent (GA) on the forgetting data, while also introducing several metrics to assess its effectiveness.
-They demonstrate that GA not only can be used for successful knowledge unlearning but also minimizes catastrophic forgetting of other knowledge within the LLMs.
-Nevertheless, GA still has significant drawbacks, such as the issue of training not converging due to unbounded loss and high unlearning cost caused by full fine-tuning for the entire LLM model.
-Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it lose knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion.
-After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other data points in the retain set ${\mathcal{D}}_{r}$ as well as preserve its previously acquired reasoning and generative capabilities.
-On the other hand, LoRA (Hu et al., 2021) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning for downstream tasks.
-The core idea behind LoRA is to freeze the pretrained LLM’s weights and train rank decomposition matrices within each layer during adaptation to downstream tasks.
-This approach not only significantly reduces the number of trainable parameters but also achieves performance comparable to full fine-tuning.
-Recently, Biderman et al.
-(2024) experimentally demonstrates that LoRA acts as a powerful regularization mechanism that prevents it from reaching the same performance levels as full fine-tuning.
-As a result, it results in less forgetting in retained knowledge.
-From this perspective, many of LoRA’s benefits are particularly advantageous in practical unlearning scenarios: 1) it significantly reduces the tuning cost for unlearning, and 2) it effectively prevents forgetting, making it a valuable method for achieving the important goal of retaining knowledge.
-Nonetheless, its application to LLM unlearning remains unexplored until now.
-In this paper, we propose a novel approach for robust and efficient unlearning for LLMs by introducing a novel loss function and leveraging LoRA with a new initialization.
-First, we analyze the derivatives of the existing GA and highlight its shortcomings: 1) dispersed gradients causing unnecessary forgetting, and 2) an unbounded loss that fails to converge.
-To address these issues, we propose the Inverted Hinge loss (IHL).
-From the derivative analysis, we show that the IHL resolves all the problems of GA, enabling faster and robust unlearning.
-Second, we explore the intersection between LLM unlearning and cost-efficient fine-tuning using LoRA.
-We find that LoRA suffers from significantly suboptimal trade-offs between its downstream performance after unlearning and its ability to generate information requested for deletion.
-To achieve robust and efficient unlearning with LoRA, we introduce a new LoRA Initialization via Weighted Low-Rank Approximation.
-Extensive experiments on the LM-Eval and Dialogue benchmarks using Neo-GPT, Phi, and Llama2 LLMs show that our proposed Inverted Hinge Loss combined with LoRA achieves robust and cost-efficient unlearning than GA (See Figure 2).
+Large Language Models (LLMs) exhibit substantial performance gains in downstream tasks with increasing model size and amount of pretraining data (Zhao et al., 2023).
+This has prompted extensive research on collecting high-quality textual corpora for LLM pretraining and developing larger models to an unprecedented scale (Brown et al., 2020; Chowdhery et al., 2023; Smith et al., 2022; Rae et al., 2021; Dubey et al., 2024).
+However, this approach has introduced significant privacy concerns due to LLMs’ tendency to memorize data indiscriminately (Carlini et al., 2021; 2023).
+For instance, Personally Identifiable Information (e.g., names, phone numbers, and email addresses) can be easily extracted from LLMs (Carlini et al., 2021).
+Additionally, OpenAI is facing multiple copyright infringement lawsuits due to unpermitted use of licensed articles during LLM pretraining (Grynbaum & Mac, 2023).
+In response to such challenges as well as increasing interest in one’s right to be forgotten (e.g., the GDPR legislation) (Voigt & Von dem Bussche, 2017; Rosen, 2011; Villaronga et al., 2018), machine unlearning for LLMs has emerged a critical and rapidly growing research field (Yao et al., 2023; Si et al., 2023).
+One method for LLM unlearning would be to filter out sensitive data from the corpus and retrain the model from scratch, an approach known as exact unlearning.
+With unprecedentedly large models and pretraining datasets, this process is highly resource-intensive and can easily become intractable under the possibility of multiple data deletion requests made in a sequential manner.
+This motivates approximate unlearning, where the goal is to remove knowledge of specific data instances without retraining the model from scratch (Figure 1).
+In this regard, Jang et al.
+(2023) proposed a simple method that finetunes LLMs using Gradient Ascent (GA) on data requested for deletion, and also introduced $n$-gram-based metrics to assess its efficacy.
+However, the proposed GA method suffers significantly not only from unstable optimization due to the objective loss being unbounded, but also from high computational cost of full-finetuning all parameters within the LLM.
+Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it forgets knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion.
+After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other relevant data points in the retain set ${\mathcal{D}}_{r}$ as well as reasoning and generative capabilities acquired from pretraining.
+Meanwhile, Low-Rank Adaptation (LoRA) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning on downstream tasks (Hu et al., 2022).
+The core idea of LoRA is to freeze all pretrained weights and instead train low-rank decomposition matrices to model the weight changes in each linear layer, effectively reducing the number of trainable parameters and thus its memory cost.
+In addition to its efficiency, the low-rankness in LoRA also induces a powerful regularization (Biderman et al., 2024), which can be beneficial in the realm of LLM unlearning by stabilizing optimization and preventing catastrophic forgetting of other remaining knowledge.
+From this perspective, we conjecture that LoRA would be a valuable approach to consider in practical unlearning scenarios, yet its application to LLM unlearning remains unexplored.
+In this paper, we explore LLM unlearning under the low-rank adaptation paradigm and propose two novel techniques for robust and efficient knowledge unlearning for LLMs.
+First, we analyze the derivatives of GA and highlight its shortcomings: 1) gradients increasing the probability of all other possible tokens cause unnecessary forgetting and 2) maximizing the next-token prediction loss involves unbounded optimization and can easily diverge.
+To address these issues, we propose the Inverted Hinge Loss (IHL) that aims to replace each token to unlearn with the next most-probable token, and show that IHL enables fast and stable tuning by resolving the issues of GA.
+Second, we find that the low-rank of LoRA imposes regularization too strong for LLM unlearning, and forcing the model to fully unlearn data points through extensive tuning leads to suboptimal cost vs.
+post-unlearning performance trade-offs.
+To accelerate unlearning in a controlled manner, we propose Fisher-Initialization of Low-rank Adapters (FILA), which data-adaptively assigns parameters responsible for generating unwanted information to adapters prior to tuning by decomposing the pretrained parameters weighted by the relative Fisher-information matrix.
+Our extensive experiments on unlearning samples in the Training Data Extraction Challenge dataset with GPT-Neo models as well as removing knowledge of fictitious author profiles using the TOFU benchmark with Phi-1.5B and Llama2-7B models consistently demonstrate that our proposed IHL combined with FILA significantly outperforms previous baselines in both efficiency and post-unlearning downstream performance.
 Figure 2: Compute cost for successful unlearning vs.
-post-unlearning downstream performances of GPT-Neo-125M.
+post-unlearning downstream performances.
+We unlearn 32 randomly sampled sequences from the Training Data Extraction Challenge from GPT-Neo-125M.
 Each point represents a different forget set and LoRA rank (if used).
@@
 Dashed lines indicate the performances of the model prior to unlearning.
-Unlearning with vanilla LoRA leads to significant loss in performance compared to full-parameter unlearning.
-However, our proposed method using both the Inverted Hinge Loss and FLoRA performs competitively to full-unlearning in all three aspects while enjoying high parameter- and cost-efficiency of LoRA.
+Unlearning via gradient differences (GD) with vanilla LoRA leads to significant loss in performance compared to full-parameter GD unlearning due to lack of plasticity.
+However, our proposed method using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the cost-efficiency of LoRA.
 2 Related Work
 Machine Unlearning.
-The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pre-trained model to discard information acquired from a specific subset of data.
-Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2021; Mehta et al., 2022; Chundawat et al., 2022; Cha et al., 2024).
-Recently, its significance has grown notably in Large Language Models (LLMs) due to its pivotal role in managing unintended memorization intrinsic to LLMs (Si et al., 2023b; Yao et al., 2024).
+The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pretrained model to discard information acquired from a specific subset of data previously used for pretraining.
+Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2023; Mehta et al., 2022; Chundawat et al., 2023; Cha et al., 2024).
+Recently, its significance has grown notably with Large Language Models (LLMs) due to crucial need for managing unintended memorization of pretraining data intrinsic to LLMs (Si et al., 2023; Yao et al., 2024b).
+Several machine unlearning algorithms have been proposed for LLMs, with most falling under the category of parameter optimization-based methods (Si et al., 2023).
 First, Wang et al.
-(2023) introduces the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets.
-Chen & Yang (2023) proposes the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model frozen.
-While these approaches demonstrates impressive unlearning results across various datasets, they are hindered by significant drawbacks, such as the necessity to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) and the need to finetune models using them (Wang et al., 2023).
-Parameter merging-based methods recently garner attention as another category of unlearning for LLMs (Wu et al., 2023; Ilharco et al., 2022).
-Experimental results demonstrate that simple task arithmetic can facilitate cost-efficient unlearning for LLMs.
-However, these methods also exhibit limitations, such as achieving a relatively weak level of unlearning (Ilharco et al., 2022) or incurring higher computational costs to detect privacy neurons (Wu et al., 2023).
-In contrast, following an unlearning approach similar to that used in image classification (Golatkar et al., 2020; Cha et al., 2024), Jang et al.
-(2023) leverages Gradient Ascent (GA) to maximize the next-token prediction loss for a sequence of tokens from the forget data.
-Unlike in image classification, they demonstrate that LLMs do not suffer from forgetting of retained knowledge when using GA on the forget data.
+(2023) introduced the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets.
+Chen & Yang (2023) proposed the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model intact.
+Liu et al.
+(2024) developed a two-stage framework: the harmful knowledge acquisition stage identifies and captures harmful knowledge within the model, and the knowledge negation stage removes that knowledge.
+While these approaches have demonstrated impressive unlearning results across various datasets, they are limited in applicability either by the need to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) or by the need to infer from a secondary model for knowledge distillation (Wang et al., 2023; Liu et al., 2024).
+Another category of LLM unlearning methods are tuning-free model-editing methods, an example of which leverages task arithmetic to suppress harmful text (Ilharco et al., 2023; Wu et al., 2023).
+While free from cost of tuning, these methods fails to show sufficient level of unlearning (Ilharco et al., 2023) or incur higher computational costs for detecting privacy neurons (Wu et al., 2023).
+In contrast, inspired by image classifier unlearning literature (Golatkar et al., 2020; Cha et al., 2024), Jang et al.
+(2023) adopted Gradient Ascent (GA) to unlearn from LLMs by maximizing the next-token prediction loss for all sequences in the forget set.
+Unlike in image classification, they demonstrate that LLMs show negligible forgetting of retained knowledge when using GA on the forget data.
 Consequently, they show that effective unlearning for LLMs is achievable solely using the forget dataset, while preserving general LLM knowledge and maintaining performance on downstream tasks.
+More recently, Yao et al.
+(2024a) demonstrated that incorporating GA with gradient descent on in-distribution data further enhances robustness against hyperparameters.
 Parameter-Efficient Fine-Tuning.
-Fine-tuning pretrained LLMs to adapt towards specific downstream tasks and instructions comes at an intractable computational cost due to the large model size and complexity.
+Full-finetuning LLMs to adapt towards specific downstream tasks and instructions incurs an intractable computational cost due to the large model size and complexity.
 To alleviate this burden, previous work has developed Parameter-Efficient Fine-Tuning (PEFT) methods that adapt only a small portion of LLM parameters while freezing the other pretrained parameters intact (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023).
-Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2021; Zhang et al., 2023; YEH et al., 2023; Kopiczko et al., 2023; Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter.
+Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2022; Zhang et al., 2023; Yeh et al., 2023; Kopiczko et al., 2024; yang Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter.
 A key benefit of LoRA is that the adapter weights can be merged seamlessly to pretrained parameters after fine-tuning, such that the post-adaptation LLM shares the same inference cost as the base pretrained LLM.
-While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposes to initialize LoRA weights using the principal singular vectors and values of the linear weights.
-Inspired by this work, we propose to extract Fisher information from the data requested for unlearning and perform weighted low-rank approximation to prepare a LoRA initialization that induces faster unlearning and better preservation of general knowledge
+While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposed to initialize LoRA weights using the principal singular vectors and values of the linear weights.
 3 Proposed Method
-In this section, we introduce the notations and provide preliminaries on unlearning through Gradient Ascent (GA) and low-rank adaptation (LoRA).
-We then analyze the gradient properties of GA and highlight its inherent issues.
-Building on this analysis, we propose the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA.
-Finally, we present the idea of employing LoRA with a novel initialization strategy, thereby enabling cost-efficient unlearning.
+In this section, we introduce useful notation and provide background information on unlearning with Gradient Ascent (GA) and Low-Rank Adaptation (LoRA).
+We then analyze derivatives of GA and highlight its inherent issues.
+Building upon this analysis, we present the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA.
+Lastly, we illustrate the method of initializing LoRA data-adaptively via Fisher-information, thereby enabling cost-efficient unlearning.
 3.1 Preliminaries
 Problem and notation.
-Given a sequence of $T$ tokens ${\bm{x}}=(x_{1},x_{2},\dots,x_{T})$, a language model (LM) models the likelihood of the sequence via next token prediction: $p_{\theta}({\bm{x}})=\prod_{t=1}^{T}p_{\theta}(x_{t}|x_{<t})$.
+Given a sequence of $T$ tokens ${\bm{x}}=(x_{1},x_{2},\dots,x_{T})$, a language model (LM) models the likelihood of the sequence via next-token prediction: $p_{\theta}({\bm{x}})=\prod_{t=1}^{T}p_{\theta}(x_{t}|x_{<t})$.
 After pretraining, we assume that an end-user has requested to delete a subset of the training set $\mathcal{D}_{f}\subset\mathcal{D}$, which we refer to as the forget set.
-Also, we consider a retain set $\mathcal{D}_{r}$ that represents the auxilary dataset which contains general knowledge the model should not forget after unlearning (e.g., Wikitext dataset (Merity et al., 2016)).
+The retain set $\mathcal{D}_{r}$ refers to an auxiliary dataset that contains other relevant knowledge that must be retained after unlearning (e.g., Wikitext; Merity et al.
+2017).
 Gradient Ascent.
-Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, which leads to a simple yet effective baseline of gradient ascent (GA) (Cha et al., 2024; Jang et al., 2023).
-Unlike the usual gradient descent, GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
-$\mathcal{L}_{GA}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$
-Note that in practice, the log-likelihood is implemented in terms of cross-entropy losses on incremental tokens.
-Thus, GA tunes the model towards maximizing the cross-entropy, the value of which is unbounded above by nature, making optimization ill-posed and unstable.
+Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, leading to a simple yet effective baseline of Gradient Ascent (GA; Jang et al.
+2023).
+GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
+$\mathcal{L}_{\text{GA}}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t})).$ (1)
+In practice, the log-likelihood is computed using cross-entropy loss and thus we GA essentially minimizes the Negative Cross-Entropy (NCE) loss.
+Therefore, GA maximizing the next-token prediction loss involves unbounded optimization, leading to an ill-posed process with unstable tuning.
+While Gradient Difference (GD) aims to alleviate this instability by minimizing the next-token prediction loss for ${\mathcal{D}}_{r}$ alongside NCE on ${\mathcal{D}}_{f}$ as regularization, we find that the approach falls short of a fundamental solution, showing performance degradation as unlearning updates are made.
 Low-Rank Adaptation.
-Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank, LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter.
-In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes
-$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}$ (1)
+Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank (Aghajanyan et al., 2021), LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter.
+In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes:
+$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}.$
 During fine-tuning, the original weight ${\bm{W}}$ is kept frozen and only the low-rank factors ${\bm{A}}$ and ${\bm{B}}$ are updated via gradient descent.
 To ensure that the initial attachment of LoRA adapters does not alter the output of the LLM, LoRA defaults to initializing ${\bm{A}}$ with a Kaiming-uniform distribution (He et al., 2015) and ${\bm{B}}$ as the zero matrix.
-Note that after fine-tuning, LoRA adapters can simply be merged with the original weights ${\bm{W}}^{\prime}={\bm{W}}+{\bm{B}}{\bm{A}}$, thereby avoiding any additional latency during inference.
-In the context of unlearning, this feature of LoRA naturally promotes a layer of security as it prevents adversarial agents from tracking previous records of unlearning just by inspecting the model architecture.
+After finetuning, LoRA adapters can simply be merged with the original weights ${\bm{W}}^{\prime}={\bm{W}}+{\bm{B}}{\bm{A}}$, thereby avoiding any additional latency during inference.
 3.2 Preliminary Results
 Despite its wide use in domain adaptation and instruction tuning, LoRA is not yet explored under the task of LLM unlearning to the best of our knowledge.
-Therefore, we first share empirical results from low-rank adapting LLMs using gradient difference (Maini et al., 2024; Liu et al., 2022a) as our objective to motivate our approach.
-For this experiment, we follow previous work (Jang et al., 2023) unlearn 32 text sequences from the Training Data Extraction Challenge from GPT-Neo-1.3B model pretrained on the Pile dataset (Gao et al., 2020).
-More information on the details on the setup can be found in Section 4 and the Appendix.
+Therefore, we first share empirical results from low-rank adapting LLMs using GD as our objective to motivate our approach.
 Figure 2 shows the results.
 Notably, vanilla LoRA suffers from lack of plasticity and ends up failing to sufficiently unlearn ${\mathcal{D}}_{f}$ within 20 epochs.
-When running more unlearning epochs or increasing the learning rate for sufficient unlearning, the model loses its previously acquired reasoning and generative capabilities, as shown in the significant decrease in LM-eval and Dialogue performances.
-In the remainder of this section, we present two techniques towards making LLM unlearning viable while enjoying the compute-efficiency of LoRA.
-3.3 Inverted Hinge Loss: Novel Loss Function for Unlearning in LLMs
-Motivation: Inherent Limitation of GA.
-We analyze the inherent issues of GA from the perspective of its gradients.
+When running more unlearning epochs or increasing the learning rate for sufficient unlearning, the model loses its previously acquired reasoning and generative capabilities, as shown in the significant decrease in Reasoning and Dialogue performances.
+In the remainder of this section, we present two techniques towards making LLM unlearning viable while enjoying the efficiency of LoRA.
+3.3 Inverted Hinge Loss: A Novel Loss Function for LLM Unlearning
+Motivation.
+We analyze the inherent issues of GA from the perspective of its derivative.
 The output layer of a language model is a softmax layer that outputs probabilities over the vocabulary.
@@
 The probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function: $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{\sum_{v=1}^{V}\exp(y_{t}^{(v)})}$ where $y_{t}^{(x_{t})}$is the logit corresponding to the true token $x_{t}$ and $y_{t}^{(v)}$ is the logit corresponding to the $v$-th token in the vocabulary.
-When we use $\mathcal{L}_{GA}$ for unlearning for LLMs, the gradient of the log-probability with respect to the logits is:
-$\frac{\partial\log(p_{\theta}(x_{t}|x_{<t}))}{\partial y_{t}^{(v)}}=\begin{cases}1-p_{\theta}(x_{t}|x_{<t})&\text{if }v=x_{t}\\ -p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
-Given this, the derivative of $\mathcal{L}_{GA}({\bm{x}})$ with respect to the logits at each $t$ becomes $\nabla_{y_{t}^{(v)}}\mathcal{L}_{GA}({\bm{x}})={\partial\log(p_{\theta}(x_{t}|x_{<t}))}/{\partial y_{t}^{(v)}}$.
-From this derivative of GA, we can confirm its unlearning mechanism and inherent issues.
-First, when $v=x_{t}$, GA decreases the prediction score corresponding to the true token while increasing the prediction score for all other tokens.
-This demonstrates that GA performs unlearning by encouraging the model to predict tokens different from the true token.
-However, despite the successful experimental results of GA (Jang et al., 2023) in LLMs, our gradient analysis reveals its inherent issues.
-When $v=x_{t}$, it increases the prediction scores of all tokens except the true token.
-As a result, GA has problems: 1) the loss not converging and becoming unbounded, and 2) unnecessary additional forgetting due to the increase in logits for all other tokens, 3) a large vocabulary size induces gradients to spread out across all other logits, making gradient updates for unlearning inefficient.
-Inverted Hinge Loss (IHL).
-Considering the limitations of GA discussed above, our motivation is to develop a simple yet novel loss function for unlearning in LLMs.
-The new loss function aims to effectively achieve unlearning by decreasing the prediction score of the true token while increasing the prediction scores of only a minimal number of other tokens.
-Inspired by the hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge Loss (IHL) as shown below:
+When we use $\mathcal{L}_{\text{GA}}$ for unlearning for LLMs, the gradient of the log-probability with respect to the logits is
+$\frac{\partial\log\left(p_{\theta}(x_{t}|x_{<t})\right)}{\partial y_{t}^{(v)}}=\begin{cases}1-p_{\theta}(x_{t}|x_{<t})&\text{if }v=x_{t}\\ -p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
+From this derivative of GA, we can interpret its unlearning mechanism: GA decreases the prediction score corresponding to the true token $x_{t}$, while increasing the prediction score for all other tokens $v\neq x_{t}$ in the vocabulary.
+Consequently, we hypothesize that GA suffers from the following problems: 1) the objective is unbounded with the possibility of divergence during optimization, 2) gradient updates are inefficient due to gradients spreading across all other logits with large vocabulary sizes, and 3) there is unnecessary degradation of generative performance due to increasing the logits for all other tokens regardless of textual coherence.
+Inverted Hinge Loss.
+To cope with aforementioned limitations of GA, we aim to design a new loss function that achieves effective unlearning by decreasing the prediction score of the true token, while focusing gradient updates on only a minimal number of viable replacements for the ground-truth token.
+Inspired by the Hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge loss (IHL) as below:
 $\mathcal{L}_{\text{IHL}}({\bm{x}})=\max\left(0,1+p_{\theta}(x_{t}|x_{<t})-\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))\right)$
-Considering the probability $p_{\theta}(x_{t}|x_{<t})$ given by the softmax function, the derivative of $\mathcal{L}_{\text{IH}}({\bm{x}})$ with respect to $y_{t}^{(v)}$ is:
+As the probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function, the derivative of $\mathcal{L}_{\text{IHL}}({\bm{x}})$ with respect to $y_{t}^{(v)}$ is:
 $\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}}.\end{cases}$
-Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IH}}({\bm{x}})\neq 0$.
-Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IH}}({\bm{x}})=0$.
-The detailed process for deriving this is provided in the Appendix A
-The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly demonstrates how the Inverted Hinge loss addresses the shortcomings of GA in knowledge unlearning for LLMs.
-First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with the opposite sign).
-This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly and intensively but also prevents spreading out of gradients for unlearning.
+Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IHL}}({\bm{x}})\neq 0$.
+Note that $\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IHL}}({\bm{x}})=0$.
+The detailed derivation can be found in Appendix A.
+The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly illustrate how the IHL addresses the shortcomings of GA in knowledge unlearning for LLMs.
+First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with opposite sign).
+This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly but also prevents spreading out of gradients.
 During this process, the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ increase slowly in proportion to the difference between the predictions for $x_{t}$ and $v^{\star}$.
 Second, once knowledge unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$), the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease.
-This not only prevents unnecessary forgetting but also results in a bounded loss form.
-3.4 FLoRA: Novel LoRA Initialization for LLM Unlearning
+This not only prevents unnecessary forgetting but also results in a bounded loss function, allowing stable unlearning via stochastic gradient descent.
+3.4 FILA: a Novel LoRA Initialization for LLM Unlearning
 Motivation.
-Aside from reduing the number of trainable parameters and thus reducing the memory requirements of LLM fine-tuning, LoRA serves as a potent form of regularization.
-While it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios (Biderman et al., 2024).
-Under the task of LLM unlearning, however, our objective strictly requires to completely remove knowledge of ${\mathcal{D}}_{f}$, and the strong stability from LoRA imposes a strong burden: vanilla LoRA unlearning requires large number of iterations through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in our Figure 2).
-Inspired by previous work on PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are important to ${\mathcal{D}}_{f}$ but not so important to ${\mathcal{D}}_{r}$ would benefit the unlearning process in terms of both convergence and post-unlearning performance.
-We hence design a novel LoRA initialization technique that measures the Fisher information on each weight ${\bm{W}}$ targeted for LoRA, and perform weighted low-rank decomposition from which we use the principal components to initialize adapter weights ${\bm{A}}$ and ${\bm{B}}$.
+Biderman et al.
+(2024) discovered that, while it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios.
+While this stability is beneficial to retaining knowledge of ${\mathcal{D}}_{r}$ during unlearning, it also imposes a strong burden under the strict objective to remove knowledge of ${\mathcal{D}}_{f}$ completely.
+As a result, vanilla LoRA unlearning requires large number of model updates through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in Figure 2).
+Drawing inspiration from PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are relatively more important to ${\mathcal{D}}_{f}$ than to ${\mathcal{D}}_{r}$ would reinforce the model’s plasticity to forgetting ${\mathcal{D}}_{f}$, thereby accelerating the rate of unlearning and minimizing catastrophic forgetting on ${\mathcal{D}}_{r}$.
+In order to quantify the importance of parameters to particular data points, we leverage Fisher information matrices.
 Parameter Importances via Fisher Information.
-The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information dataset ${\mathcal{D}}$ provides on model parameters $\theta$.
+The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information a dataset ${\mathcal{D}}$ provides on model parameters $\theta$.
 More concretely, ${\bm{F}}_{\theta}$ is computed as the second cross-moments of first partial derivatives of the log-likelihood of ${\mathcal{D}}$ (left of Eq.
-3.4).
-However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead which, in the context of LLMs, can be computed as the average of squared gradients propagated from the language modeling objective (i.e.
-cross entropy loss):
-${\bm{F}}_{\theta}=\mathbb{E}_{\mathcal{D}}\left[\left(\dfrac{\partial}{\partial\theta}\log p{\color[rgb]{0,0,0}{}_{\theta}}({\mathcal{D}}|\theta)\right)^{2}\right]\approx\dfrac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}\in{\mathcal{D}}}\left(\dfrac{\partial}{\partial\theta}\mathcal{L}_{\text{LM}}({\bm{x}};\theta)\right)^{2}\eqcolon\hat{{\bm{F}}}_{\theta},$ (2)
-where $\mathcal{L}_{\text{LM}}$ is the next-token prediction loss of LM, such as $\mathcal{L}_{\text{LM}}({\bm{x}};\theta)=\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$.
-Within our LLM unlearning setup, a high empirical Fisher information measured with ${\mathcal{D}}_{f}$ indicates that the next-token prediction loss on ${\mathcal{D}}_{f}$ leads to large absolute gradients on the parameter under concern, and we consider such parameters to be important in generating sequences in ${\mathcal{D}}_{f}$.
-Let $\hat{{\bm{F}}}_{\bm{W}}^{f}$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp.
+2).
+However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead.
+In the context of LLMs, this can be computed as:
+${\bm{F}}_{\theta}({\mathcal{D}})=\mathbb{E}_{\mathcal{D}}\left[\left(\dfrac{\partial}{\partial\theta}\log p{{}_{\theta}}({\mathcal{D}}|\theta)\right)^{2}\right]\approx\dfrac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}\in{\mathcal{D}}}\left(\dfrac{\partial}{\partial\theta}\mathcal{L}_{\text{LM}}({\bm{x}};\theta)\right)^{2}\eqcolon\hat{{\bm{F}}}_{\theta}({\mathcal{D}}),$ (2)
+where $\mathcal{L}_{\text{LM}}$ denotes the next-token prediction loss used to pretrain LMs, $\mathcal{L}_{\text{LM}}({\bm{x}};\theta)=\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$.
+Within our LLM unlearning setup, a high empirical Fisher information measured with ${\mathcal{D}}_{f}$ indicates that $\mathcal{L}_{\text{LM}}$ on ${\mathcal{D}}_{f}$ leads to large absolute gradients on the parameter under concern, and we consider such parameters to be important in generating sequences in ${\mathcal{D}}_{f}$.
+Let $\hat{{\bm{F}}}_{\bm{W}}^{f}\coloneqq\hat{{\bm{F}}}_{\bm{W}}({\mathcal{D}}_{f})$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp.
 $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$).
-Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{rel}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to first identify parameters that are important specifically for ${\mathcal{D}}_{f}$, but unimportant to ${\mathcal{D}}_{r}$.
-While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus choose to use the relative Fisher information matrix rather than $\hat{{\bm{F}}}^{f}$.
-Fisher-weighted LoRA Initialization (FLoRA).
-Intuitively, initializing LoRA weights to be parameters that are important in generating ${\mathcal{D}}_{f}$ would benefit unlearning as gradients would only modify parameters important to ${\mathcal{D}}_{f}$ while keeping all other parameters intact.
-Given the relative importance $\hat{{\bm{F}}}_{\bm{W}}^{rel}$ for each target parameter ${\bm{W}}$, we propose to initialize LoRA adapter weights as the solution to the following problem of Weighted Low-Rank Approximation (WLRA):
-$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\sum_{i,j}\left([\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}({\bm{W}}-{\bm{B}}{\bm{A}})_{i,j}\right)^{2}$ (3)
-Note that when all weights equal one, this problem reduces down to a standard low-rank approximation problem, of which the solution can easily be computed via rank-$r$ SVD.
-In general, however, this minimization problem does not have a closed-form solution and thus requires iterative optimization (Srebro & Jaakkola, 2003).
-While we may resort to iterative methods towards initializing LoRA weights in a fine-grained manner, this would undermine the gains in efficiency due to employing LoRA.
-Thus, we tackle a simplified version of WLRA instead by assuming that parameters in each row of ${\bm{W}}$ share the same importance, for which we use the square-root of the row-wise sum of the relative Fisher information matrix as
-$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\left\|\text{diag}\left((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\ \bm{\cdot}\ \mathds{1})^{\frac{1}{2}}\right)({\bm{W}}-{\bm{B}}{\bm{A}})\right\|_{2}$ (4)
-where $\mathds{1}\in\mathbb{R}^{k}$ is the all-one vector.
-$\text{diag}(\cdot)$ and $\bm{\cdot}$ denote the vector diagonalization function and the matrix-vector product, respectively.
-Unlike general WLRA, the row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
-Given the solution, we use the solution ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights.
-To ensure that the output is not perturbed after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}{\bm{A}}$.
-In summary, our Fisher-Weighted initialization extracts parameters that are important only to ${\mathcal{D}}_{f}$ and not to ${\mathcal{D}}_{r}$ such that LoRA tuning can be focused towards erasing knowledge of ${\mathcal{D}}_{f}$.
-3.5 Final Loss Function for Unlearning
-Following initialization with FLoRA, unlearning is performed on the model $\Theta=\theta\cup\theta_{FLoRA}$, which consists of the original weights $\theta$ and the LoRA weights for each layer $\theta_{FLoRA}=\{{\bm{A}}^{*}_{l},{\bm{B}}^{*}_{l}\}_{l=1}^{L}$, where $L$ represents the number of layers where FLoRA is applied.
-During unlearning process, $\theta$ remains frozen, and only updates to the FLoRA weights $\theta_{FLoRA}$ are made.
-Furthermore, when performing unlearning on the forget dataset $D_{f}$, we incorporate the auxiliary retain set $D_{r}$, which contains general knowledge (e.g., the Wikitext dataset), as a default setting, following the gradient difference approach (Liu et al., 2022a; Maini et al., 2024).
-The final loss function using the proposed Inverted Hinge Loss is defined as follows:
-$\operatorname*{minimize}_{\theta_{FLoRA}}\sum_{{\bm{x}}_{r}\in D_{f},{\bm{x}}_{f}\in D_{r}}\mathcal{L}_{\text{IH}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (5)
-In practice, training (unlearning) for the LLM model involves sampling corpora from each dataset, $D_{f}$ and $D_{r}$, and minimizing the final loss function through stochastic gradient descent.
+Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to identify parameters that are important exclusively for ${\mathcal{D}}_{f}$ and not for ${\mathcal{D}}_{r}$.
+While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus use $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$ rather than $\hat{{\bm{F}}}_{\bm{W}}^{f}$.
+Fisher-weighted Initialization of Low-rank Adapters.
+Given the relative importance $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$ for each target weight ${\bm{W}}$, we propose to initialize the corresponding LoRA adapter weights with the solution to the following Weighted Low-Rank Approximation (WLRA) problem:
+$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\sum_{i,j}\left([\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}({\bm{W}}-{\bm{B}}{\bm{A}})_{i,j}\right)^{2}.$
+Note that when all weights $[\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}$ equal one, WLRA reduces to standard low-rank matrix approximation, for which the solution can easily be computed via rank-$r$ SVD.
+For general weights, however, this minimization problem does not have a closed-form solution and requires iterative optimization (Srebro & Jaakkola, 2003).
+While we may resort to iterative methods to initialize LoRA weights in a fine-grained manner, this would undermine the efficiency gains from deploying low-rank adapters.
+Therefore, we assume that parameters in each row of ${\bm{W}}$ share the same importance equal to the square-root of the row-wise sum of $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$, and simplify the problem to
+$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\left\|\texttt{diag}\left((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{\frac{1}{2}}\right)({\bm{W}}-{\bm{B}}{\bm{A}})\right\|_{2}$
+with $\mathds{1}\in\mathbb{R}^{k}$ and $\texttt{diag}(\cdot)$ indicating the all-one vector and the vector diagonalization function, respectively.
+Unlike general WLRA, this row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
+Given this solution, we use ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights.
+To ensure that the model behavior remains the same after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}^{*}{\bm{A}}^{*}$.
+Intuitively, our Fisher-weighted Initialization of Low-rank Adapters (FILA) extracts parameters that are important for generating ${\mathcal{D}}_{f}$, but not for generating ${\mathcal{D}}_{r}$, such that LoRA tuning can be focused on erasing knowledge relevant to ${\mathcal{D}}_{f}$ while keeping information regarding ${\mathcal{D}}_{r}$.
+3.5 Final Loss Function for LLM Unlearning
+In summary, we perform unlearning on the model $\bm{\Theta}=\bm{\theta}\cup\bm{\theta}_{\text{FILA}}$, consisted of original pretrained weights $\bm{\theta}$ and the FILA-initialized low-rank adapter weights for each linear layer $\bm{\theta}_{\text{FILA}}=\{{\bm{A}}^{*}_{\ell},{\bm{B}}^{*}_{\ell}\}_{\ell=1}^{L}$, where $L$ represents the number of layers tuned via LoRA.
+Additionally, we incorporate GD, which utilizes the auxiliary retain set $\mathcal{D}_{r}$.
+The final loss function using both the proposed IHL and FILA is defined as follows:
+$\operatorname*{minimize}_{\theta_{\text{FILA}}}\sum_{{\bm{x}}_{r}\in{\mathcal{D}}_{f},{\bm{x}}_{f}\in{\mathcal{D}}_{r}}\mathcal{L}_{\text{IHL}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (3)
+In practice, training (unlearning) for the LLM model is conducted by minimizing Eq.3 through stochastic gradient descent.
 4 Experiments
-In this section, we first perform unlearning experiments based on the Training Data Extraction Challenge (§4.1), followed by ablation and analytical results (§4.2).
-All experiments were conducted on a remote GCP server equipped with 16 NVIDIA A100 Tensor Core GPUs.
+In this section, we first perform experiments unlearning samples from the Training Data Extraction Challenge (TDEC; §4.1), followed by ablation and analytical results (§4.2).
+We also conduct experiments on the Task of Fictitious Unlearning (TOFU; §4.3), a benchmark that well-mimics a real-world scenario for LLM unlearning evaluation.
+For brevity, we present results from additional experiments such as continual unlearning in Appendix D.
 Table 1: Evaluation results on LM-Evaluation, Dialogue datasets before and after unlearning samples from the TDEC dataset.
-Note that, in all experiments using LoRA, GD is used as the default loss function.
-The "+IHL" results refer to experiments where GA within GD is replaced with the proposed IHL (i.e., using Eqn.(5)), and "+FLoRA" indicates the results where FLoRA is used instead of LoRA.
+For all runs under LoRA, we use rank 16 and GD is used as the baseline.
+The "+IHL" results refer to experiments where NCE loss within GD is replaced with our proposed IHL (i.e., Equation 3), and "+FILA" indicates results where FILA is used to initialize LoRA.
 Model Method Params. Epochs EL10 MA Reasoning Dialogue Pile
@@
 + IHL 11.4 0.4 22.7 41.9 6.0 32.9
-+ FLoRA 6.0 0.3 23.9 42.2 10.1 24.0
+     + FILA 6.0 0.3 23.9 42.2 10.1 24.0
 Before - - 67.6 92.2 49.8 11.5 11.5
@@
 + IHL 20.0 1.7 44.6 47.1 10.2 14.9
-+ FLoRA 13.0 0.5 29.6 48.3 12.1 14.7
+     + FILA 13.0 0.5 29.6 48.3 12.1 14.7
 Before - - 70.4 93.4 52.3 11.5 10.4
@@
 + IHL 17.8 0.0 26.7 49.6 8.5 22.2
-+ FLoRA 10.3 0.1 28.5 49.6 10.7 16.0
+     + FILA 10.3 0.1 28.5 49.6 10.7 16.0
 4.1 Training Data Extraction Challenge
 Experimental Setup.
-Assuming a scenario where an adversary attacks a pretrained LLM to extract text samples previously used in training, the Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) that are found to be easily extractable.
+The Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) found to be easily extractable from a pretrained LLM.
 For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$.
-For the retain set ${\mathcal{D}}_{r}$, we use the validation subset of WikiText as it contains factual world knowledge that we wish to maintain within the LLM.
-For TDEC experiments, we consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using three different forget sets.
-For this experiment, we use a fixed learning rate of 5e-5 when full finetuning, and 2e-4 when using LoRA and FLoRA.
-For all model sizes, we use a LoRA (and FLoRA) rank of $r=16$ as a default.
-We set the default to applying LoRA to the attention matrices for Query (Q) and Value (V), as well as to the Feedforward Network (FFN).
-Following previous work (Jang et al., 2023), we measure the unlearning efficacy via two metrics.
+For the retain set ${\mathcal{D}}_{r}$, we use the subset of WikiText (Merity et al., 2017) as it contains factual world knowledge that we wish to maintain after unlearning.
+We consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using five different forget sets.
+For this experiment, we use a fixed learning rate of 2e-4 and use LoRA adapters with rank $r=\{4,8,16,32\}$.
+For reasons we illustrate later in §4.2, we choose to apply LoRA on query and value layers in the attention module and two linear layers within feed-forward layers.
+Following previous work (Jang et al., 2023), we measure the unlearning efficacy with two metrics.
 The ${\bm{n}}$-gram Extraction Likelihood (ELn) measures the $n$-gram overlap between the ground truth sequence in ${\mathcal{D}}_{f}$ and the output generated by the model.
 The Memorization Accuracy (MA) measures the token-wise accuracy of the LLM on ${\mathcal{D}}_{f}$.
-More details on these metrics are introduced in Section B of the Appendix.
-After each unlearning epoch, we measure EL10/AM and consider the model to have successfully unlearned ${\mathcal{D}}_{f}$ if both values are smaller than those measured using a held-out validation set.
-Once unlearning is finished, we evaluate the unlearned model on various downstream benchmarks to ensure the LLM maintains its previously acquired language modeling capabilities after unlearning.
-To assess its reasoning capabilities, we average accuracies across 9 different classification datasets LAMBADA (Paperno et al., 2016), Hellaswag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), COPA (Gordon et al., 2012), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), PiQA (Bisk et al., 2020), MathQA (Amini et al., 2019), PubmedQA (Jin et al., 2019).
-To measure generative performance, we also measure the F1 score averaged over four dialogue generation datasets, Blended Skill Talk (Smith et al., 2020), Empathetic Dialogues Rashkin et al.
-(2019), Wizard of Internet (Komeili et al., 2021), Wizard of Wikipedia (Dinan et al., 2018).
+More details on these metrics are shared in Appendix B.
+After each unlearning epoch, we measure EL10 and MA of the model, and we consider the model has successfully unlearned ${\mathcal{D}}_{f}$ if both values measured on ${\mathcal{D}}_{f}$ become smaller than those measured from a held-out validation set that the model has never seen before within 20 unlearning epochs.
+Once unlearning is finished, we evaluate the unlearned model on various downstream benchmarks to measure how well the LLM maintains its previously acquired reasoning and generative capabilities.
+To assess its reasoning capabilities, we average accuracies across 9 different classification datasets.
+To measure generative performance, we also average the F1 scores over four dialogue generation datasets.
 Lastly, we measure the perplexity on the validation subset of the Pile (Gao et al., 2020).
-Baseline.
-For experiments on knowledge unlearning in LLMs, we select using two baseline algorithms, Gradient Ascent (GA) (Jang et al., 2023) and Gradient Difference (GD) (Liu et al., 2022a; Maini et al., 2024), which rely solely on the original language model while utilizing a forget dataset $D_{f}$ only or both a forget dataset $D_{f}$ and an auxiliary dataset $D_{r}$, respectively.
-For all experiments, we use the same forgetting criteria for ELn and MA as in  Jang et al.
-(2023) and report the results after performing unlearning until these criteria are met.
-Figure 3: Results from unlearning 32 examples from Training Data Extraction Challenge.
-Each row represents the average performance in each set of LM capabilies: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better).
-Each column shows results under different model size: 125M, 1.3B, and 2.7B.
-Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed.
-X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics.
+A comprehensive list of evaluation datasets can be found in Appendix C.
+We consider two LLM unlearning baselines, Gradient Ascent (GA) (Jang et al., 2023) and Gradient Difference (GD) (Liu et al., 2022a; Maini et al., 2024), both of which only require the original language model and datasets ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$ representing knowledge we wish to unlearn and retain, respectively.
+We exclude methods that require another auxiliary model (Wang et al., 2023; Liu et al., 2024) or the entire training data (Wang et al., 2023; Chen & Yang, 2023) from our baselines.
+Figure 3: Results from unlearning examples in the TDEC dataset from the GPT-Neo LLMs with three different sizes.
+Each row represents the performance averaged across datasets within each set of LLM capability tests: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better).
+The circles and crosses represent successful and unsuccessful attempts, respectively, of unlearning a particular forget set ${\mathcal{D}}_{f}$.
 Solid lines indicate the performance of different methods averaged only across successful unlearning trials.
-The dashed lines show respective performances prior to unlearning.
-Unlearning with vanilla GD not only leads to significant loss in performance, but also fails to unlearn sufficiently even with large LoRA ranks.
-Replacing the negative cross-entropy loss with our hinge loss (IHL) shows significant gains in reasoning and generation capabilities, but still fails to unlearn a number of example sets.
-Using NCE but with Fisher-weighted LoRA initialization (FLoRA) notably increases the rate of successful unlearning, but at the cost of significant loss in overall LM capability.
-Using both IHL and FLoRA best minimizes post-unlearning performance degradation in all three capabilities under successful unlearning.
+The dashed lines indicate the base model performance prior to unlearning.
+Unlearning with GD leads to significant loss in performance and also frequently fails to unlearn effectively even with large LoRA ranks.
+Replacing the negative cross-entropy (NCE) loss with our hinge loss (IHL) boosts retention of reasoning and generation capabilities, but still fails to unlearn in multiple cases.
+Using NCE loss with FILA initialization notably increases the rate of unlearning success, but at the cost of losing overall performance.
+Using both IHL and FILA best minimizes post-unlearning performance degradation in all three aspects.
 Results.
-Table 1 shows the evaluation results of full-parameter unlearning and unlearning with LoRa and FLoRA.
-From our experimental results, we can identify the following findings.
-First, for all model sizes, using GD for unlearning not only achieves the forgetting criteria in fewer epochs compared to using GA alone, but it also better preserves previously acquired knowledge.
-Specifically, for the 125M and 1.3B models, knowledge unlearning with GA leads to a significant deterioration in PPL on the Pile dataset, whereas GD not only overcomes this issue but also achieves improved Dialogue (F1) performance compared to Before (e.g., GPT-Neo 1.3B).
-Second, we observe that simply applying GD loss with LoRA fails to achieve successful unlearning across all model sizes.
-Despite the fact that only about 0.7% to 1.6% of the total parameters are updated for unlearning when using LoRA, its simple application actually causes forgetting of previously acquired knowledge.
-Finally, when using the proposed IHL along with FLoRA, it achieves the forgetting criteria in fewer epochs while better preserving previously acquired knowledge.
-Although using IHL with LoRA alone maintains previously acquired knowledge better than the baseline, it requires relatively more epochs to meet the forgetting criteria.
-However, when combined with FLoRA, not only is the number of epochs required to meet the forgetting criteria significantly reduced, but overall, previously acquired knowledge is preserved more successfully.
-Additionally, we conduct further experiments with different ranks for LoRA and FLoRA, with results and detailed explanations provided in Figure 3 and its caption.
-From all experiments so far, we could confirm that simply applying LoRA to GD does not lead to successful unlearning.
-However, using the proposed IHL in combination with FLoRA demonstrates better cost-efficiency than the existing GA and GD methods, while also better preserving previously acquired knowledge during unlearning.
-Figure 4: Results from unlearning 32 examples from Training Data Extraction Challenge using LoRA for GPT-Neo 2.7B.
-Note that each color denotes Unlearning with vanilla GD, Replacing the negative cross-entropy loss with our hinge loss (IHL), Using NCE but with Fisher-weighted LoRA initialization (FLoRA), and Using both IHL and FLoRA.
-Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed.
-X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics.
+Table 1 shows evaluation results using a fixed LoRA rank of 16 and Figure 3 shows analogous results using a different LoRA ranks.
+Our key findings are as follows.
+First, using GD not only meets the forgetting criteria in fewer epochs across all model sizes but it also preserves previously acquired knowledge (i.e., performance on Reasoning, Dialogue, and Pile) better than GA.
+Mainly for 125M and 1.3B models, GA causes a significant decline in generative performance, whereas GD partially mitigates this decline and improves both Dialogue F1 scores and the Pile perplexity.
+Second, we find that simply applying GD with LoRA fails to achieve successful unlearning across all model sizes.
+Second, we find that imply running GD with LoRA significantly increases the number of epochs required for successful unlearning.
+While enjoying great parameter-efficiency by tuning only about 0.7% to 1.6% of the total parameters when using LoRA, its application to unlearning with GD results in significant loss in overall language capability, especially on generative tasks.
+Lastly, using our proposed IHL in conjunction with FILA achieves the forgetting criteria in fewest epochs while best preserving LLM performance.
+Although IHL with LoRA alone retains knowledge better than the baseline, it requires more epochs to meet the forgetting criteria.
+This is resolved when used together with FILA, and the number of required epochs is significantly reduced, reducing catastrophic forggetting of previously acquired knowledge even further.
+Figure 4: Results from unlearning examples from TDEC dataset using LoRA with rank 32 to adapt sets of layers on GPT-Neo 2.7B.
+The marker shapes and colors are used similarly as in Figure 3.
+Based on the rate of unlearning success, tuning FFN layers (e.g., FFN, QVFFN) is more receptive to targeted knowledge removal compared to tuning attention layers (e.g., QV, QKVO).
 4.2 Analysis
 What modules do we need to adapt?
-Figure 4 presents experiments where LoRA or FLoRA is applied to various weight matrices, including those for Query (Q), Value (V), Key (K), Output (O) in the self-attention module, and the FeedForward Network (FFN).
-While the original LoRA paper indicates that applying LoRA to Q and V yields superior performance on downstream tasks (Hu et al., 2021), our experiments show that applying LoRA to Q and V alone is insufficient to meet the unlearning criteria in all cases.
-Notably, applying LoRA to the FFN results in much more successful unlearning in cases of using FLoRA.
-Furthermore, applying FLoRA to QVFFN in conjunction with IHL not only achieves superior unlearning results but also better preserves previously acquired knowledge.
-Based on these findings, we adopt the application of LoRA and FLoRA to QVFFN as the default setting.
+Figure 4 presents experiments where low-rank adapters are attached to various target parameter groups, including those for Query (Q), Value (V), Key (K), Output (O) in the attention module, and the Feed-Forward Network (FFN).
+While the original LoRA paper (Hu et al., 2022) indicates that applying LoRA to Q and V yields superior performance on downstream tasks, our experiments indicate that using LoRA on Q and V only is insufficient to meet the unlearning criteria within our timeframe of 20 epochs.
+Notably, when LoRA is applied to FFNs, we observe significant increase in rate of successful unlearning.
+Furthermore, integrating FILA with IHL achieves the best post-unlearning performance across all LoRA target module combinations.
 Cost-efficiency of the proposed method.
-To compare unlearning costs, we evaluate the FLOPs for each method and the results are presented in Figure 2.
-While using vanilla LoRA offers significant advantages in terms of unlearning cost (i.e., FLOPs), as confirmed by our previous experiments, it results in substantial performance losses compared to full-parameter unlearning.
-In contrast, combining the proposed Inverted Hinge Loss with FLoRA not only achieves the best performance but also leverages the cost advantages of LoRA, demonstrating that this approach enables cost-efficient unlearning.
+Our compute-cost vs.
+performance comparisons in Figure 2 show that, while vanilla LoRA allows significant reduction in unlearning costs (i.e., FLOPs) by freezing the majority of parameters, it incurs substantial performance losses compared to full-parameter unlearning due to excessive stability originating from its low-rankness.
+In contrast, combining the proposed IHL with FILA not only achieves the best performance but also leverages the cost advantages of LoRA.
+4.3 Task of Fictitious Unlearning
+Experimental Setup.
+The Task of Fictitious Unlearning (TOFU) benchmark (Maini et al., 2024) is a synthetic dataset containing 20 question-answer pairs for each of 200 fictitious author profiles generated by GPT-4.
+The TOFU evaluation pipeline first finetunes a pretrained LLM on all QA pairs.
+Given this finetuned LLM that serves as our base model, our task is to unlearn all information regarding 1%, 5%, or 10% of the authors from the model.
+Note that we can obtain reference models finetuned only on the retain set (QA-pairs on 99%, 95%, or 90% of authors), with which we evaluate the Forget Quality of unlearned models by measuring the p-value from a Kolmogorov-Smirnov test.
+A high p-value indicates high distributional similarity between the unlearned model and the reference model, thus implying strong forgetting.
+To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregated model performance.
+Further details on the dataset and evaluation pipeline can be found in  Maini et al.
+(2024).
+Following the original paper of TOFU, we prepare two base models by finetuning Phi-1.5B and Llama2-7B on TOFU for 5 epochs with learning rates 2e-5 and 1e-5, respectively.
+We then unlearn using two baselines (GA and GD) and our two methods (IHL and IHL+FILA) using LoRA adapters of rank 4, 8, 16, or 32.
+For unlearning, we use a learning rate of 2e-4 if our base model is from Phi-1.5B and 1e-4 for Llama2-7B.
+All training procedures run 5 epochs with an effective batch size of 32 using the AdamW optimizer (Loshchilov & Hutter, 2019).
+(a) Phi-1.5B
+(b) Llama2-7B
+[…]
+Results.
+Figure 5 shows the model utility vs.
+forget quality curves from unlearning three differently-sized TOFU forget sets from Phi-1.5B and Llama2-7B models.
+Comparing results among different forget set sizes, we first observe that forgetting 1% of author profiles is fairly straightforward, as all curves quickly approach the reference model with a single epoch, with increasing the LoRA rank leading to incremental improvements in performance.
+On the other hand, when unlearning a larger set of profiles (i.e., 5% or 10%), we see that both GA and GD quickly degrades model utility.
+With regards to our proposed method, we find that replacing the NCE loss in GD with our IHL better retains model utility across all LoRA ranks and forget set sizes, as curves are more aligned straight-up towards the reference point with negligible shift in model utility.
+This stability comes at the cost of unlearning efficiency, however, as the rate at which the LLM forgets ${\mathcal{D}}_{f}$ is slower with IHL due to IHL decreasing the likelihood of the unwanted token by increasing the likelihood of only one other most-possible token each time in a controlled manner.
+Nonetheless, initializing LoRA adapters with FILA largely alleviates this issue and enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.
+Interestingly, we find the prior weight assignment via FILA can lead to excessive unlearning in some cases (e.g., unlearning 10% forget set with ranks 8 or 16 on Llama2-7B), with model updates reducing the forget quality after reaching the upper bound at zero.
+This behavior resembles the Streisand effect as unlearning gradients beyond a certain point in optimization unintentionally renders ${\mathcal{D}}_{f}$ more noticeable within the model (Golatkar et al., 2020).
+As reference models are not available for measuring forget quality in real-world scenarios, finding the optimal point at which to stop unlearning to prevent this effect as well as designing a robust evaluation metric that does not depend upon oracle models would be interesting directions, which we leave as future work.
 5 Concluding Remarks
-In this paper, we address limitations of existing unlearning techniques (i.e., Gradient Ascent) by introducing the novel Inverted Hinge Loss (IHL) and leveraging LoRA with a novel initialization (i.e., Fisher-weighted LoRA Initialization (FLoRA)).
-Our findings reveal that IHL overcomes key drawbacks associated with Gradient Ascent (GA), such as dispersed gradients and unbounded loss, leading to faster and more robust unlearning.
-Furthermore, our approach of combining FLoRA with IHL not only reduces the cost of unlearning but also enhances the preservation of previously acquired knowledge.
-Through experiments, we show that our method achieves superior performance in knowledge unlearning compared to the baseline method.
+In this paper, we address limitations of Gradient Ascent (GA), a widely used method for LLM unlearning, and introduce a novel Inverted Hinge loss (IHL) to replace the negative cross-entropy loss in GA and resolve issues with dispersed gradients and unboundedness.
+We also propose Fisher-weighted initialization for low-rank adaptation (FILA) that pre-assigns weights relatively important to generating unwanted information as means to facilitate efficient LLM unlearning with LoRA.
+Experiments on the Training Data Extraction Challenge dataset with GPT-Neo models along with the TOFU benchmark using Phi-1.5B and Llama2-7B models show that our proposed methods enable faster and more stable LoRA-based LLM unlearning, significantly outperforming existing baselines in computational efficiency as well as post-unlearning performance.
 Appendix A Derivative Analysis for the Inverted Hinge Loss Function
@@
 $\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(x_{t})}}=\frac{\exp(y_{t}^{(x_{t})})}{Z}-\frac{\exp(y_{t}^{(x_{t})})^{2}}{Z^{2}}=p_{\theta}(x_{t}|x_{<t})-p_{\theta}(x_{t}|x_{<t})^{2}=p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))$
-For the first case, using the chain rule again, we get:
+For the second case, using the chain rule again, we get:
 $\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(k)}}=\frac{\partial}{\partial y_{t}^{(k)}}\left(\frac{\exp(y_{t}^{(x_{t})})}{Z}\right)=-\frac{\exp(y_{t}^{(x_{t})})}{Z^{2}}\frac{\partial Z}{\partial y_{t}^{(k)}}$
@@
 $\dfrac{1}{\mathcal{D}_{f}}\sum_{{\bm{x}}\in\mathcal{D}_{f}}\textsc{EL}_{10}({\bm{x}})\leq\dfrac{1}{\mathcal{D}_{\text{val}}}\sum_{{\bm{x}}\in\mathcal{D}_{\text{val}}}\textsc{EL}_{10}({\bm{x}})\;\;\;\;\text{and}\;\;\;\;\dfrac{1}{\mathcal{D}_{f}}\sum_{{\bm{x}}\in\mathcal{D}_{f}}\textsc{MA}({\bm{x}})\leq\dfrac{1}{\mathcal{D}_{\text{val}}}\sum_{{\bm{x}}\in\mathcal{D}_{\text{val}}}\textsc{MA}({\bm{x}})$
+Appendix C Additional Details on Experimental Setting
+Experiemtal Settings   All experiments were conducted on a remote server equipped with NVIDIA A100 40GB Tensor Core GPUs.
+Datasets for Evaluation in the TDEC   To evaluate reasoning capabilities, we utilize nine different classification datasets: LAMBADA (Paperno et al., 2016), Hellaswag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), COPA (Gordon et al., 2012), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), PiQA (Bisk et al., 2020), MathQA (Amini et al., 2019), and PubmedQA (Jin et al., 2019).
+To assess generative performance, we employ Blended Skill Talk (Smith et al., 2020), Empathetic Dialogues (Rashkin et al., 2019), Wizard of Internet (Komeili et al., 2022), and Wizard of Wikipedia (Dinan et al., 2019).
+Details of metrics of TOFU   We evaluate the Forget Quality of unlearned models by measuring the p-value from the Kolmogorov-Smirnov test that compares the empirical distribution of our unlearned model to that of the reference model.
+To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregated model performance on the retain set of remaining fictitious author profiles, and two held-out sets consisted of QA-pairs regarding real author profiles and other world facts.
+Appendix D Additional Experimental Results
+D.1 Continual Unlearning
+Because of the importance of continual unlearning (or sequential unlearning) in real-world applications, previous studies have underscored its relevance through a sequence of unlearning tasks (Cha et al., 2024; Jang et al., 2023).
+Building on them, we conduct continual unlearning experiments involving four tasks.
+Figure 6 of the Appendix shows that IHL consistently outperforms GD across all metrics.
+Notably, the proposed IHL demonstrates significantly enhanced performance on the four Dialogue and Pile datasets.
+Finally, we confirm that the combination of IHL and FLoRA achieves more robust and cose-efficient continual unlearning, as evidenced by the experimental results for Reasoning, Dialogue, and Pile, while utilizing only about 1.6% of the total parameters.
+Figure 6: Experimental results of continual unlearning.
+Each task consists of 32 disjoint sequences sampled from the TDEC dataset, leading to a total of 128 sequences to unlearn.
+For these experiments, we use the pretrained GPT-Neo 125M model.
+The experimental setup for unlearning and the forgetting criteria are configured as in the previous TDEC experiments.
+Task 0 refers to the result before unlearning.

### v3

@@
-Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models
+Towards Robust and Parameter-Efficient Knowledge Unlearning for LLMs
 Abstract
@@
 While Gradient Ascent (GA) is commonly used to unlearn by reducing the likelihood of generating unwanted content, it leads to unstable optimization and catastrophic forgetting of retrained knowledge.
-We also find that combining GA with low-rank adaptation results in poor trade-offs between computational cost and generative performance.
+We find that combining GA with low-rank adaptation results in poor trade-offs between computational cost and generative performance.
 To address these challenges, we propose two novel techniques for robust and efficient unlearning for LLMs.
@@
 Experiments on the Training Data Extraction Challenge dataset using GPT-Neo models as well as on the TOFU benchmark with Phi-1.5B and Llama2-7B models demonstrate that our approach effectively removes sensitive information while maintaining reasoning and generative capabilities with minimal impact.
+Our implementation can be found in https://github.com/csm9493/efficient-llm-unlearning.
 1 Introduction
@@
 This motivates approximate unlearning, where the goal is to remove knowledge of specific data instances without retraining the model from scratch (Figure 1).
-In this regard, Jang et al.
-(2023) proposed a simple method that finetunes LLMs using Gradient Ascent (GA) on data requested for deletion, and also introduced $n$-gram-based metrics to assess its efficacy.
-However, the proposed GA method suffers significantly not only from unstable optimization due to the objective loss being unbounded, but also from high computational cost of full-finetuning all parameters within the LLM.
-Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it forgets knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion.
-After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other relevant data points in the retain set ${\mathcal{D}}_{r}$ as well as reasoning and generative capabilities acquired from pretraining.
+In this regard, several novel approaches have been proposed for approximate unlearning: Jang et al.
+(2023) introduced a simple method that finetunes LLMs using Gradient Ascent (GA) on data requested for deletion and also proposed $n$-gram-based metrics to evaluate its effectiveness.
+Wang et al.
+(2023) and Liu et al.
+(2024) proposed knowledge distillation-based methods that selectively transfer knowledge to a secondary model for unlearning.
+However, both approaches face significant challenges: GA suffers from unstable optimization due to unbounded nature of its objective loss, while distillation-based methods incur substantial computational costs from relying on a secondary model.
+Above all, these approaches share a critical drawback: the high computational cost of full fine-tuning all parameters within the LLMs.
+Figure 1: LLM unlearning aims to forget data points in ${\mathcal{D}}_{f}$ while maintaining knowledge of the retain set ${\mathcal{D}}_{r}$.
+Unlike GA, our IHL induces higher unlearning stability by reducing the likelihood of unwanted tokens in a controlled manner.
+To accelerate unlearning with IHL, FILA extracts and places parameters important in generating $\mathcal{D}_{f}$ to LoRA weights a priori via weighted low-rank approximation.
+IHL and FILA form a powerful synergy towards robust and efficient LLM unlearning.
 Meanwhile, Low-Rank Adaptation (LoRA) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning on downstream tasks (Hu et al., 2022).
 The core idea of LoRA is to freeze all pretrained weights and instead train low-rank decomposition matrices to model the weight changes in each linear layer, effectively reducing the number of trainable parameters and thus its memory cost.
-In addition to its efficiency, the low-rankness in LoRA also induces a powerful regularization (Biderman et al., 2024), which can be beneficial in the realm of LLM unlearning by stabilizing optimization and preventing catastrophic forgetting of other remaining knowledge.
-From this perspective, we conjecture that LoRA would be a valuable approach to consider in practical unlearning scenarios, yet its application to LLM unlearning remains unexplored.
-In this paper, we explore LLM unlearning under the low-rank adaptation paradigm and propose two novel techniques for robust and efficient knowledge unlearning for LLMs.
-First, we analyze the derivatives of GA and highlight its shortcomings: 1) gradients increasing the probability of all other possible tokens cause unnecessary forgetting and 2) maximizing the next-token prediction loss involves unbounded optimization and can easily diverge.
-To address these issues, we propose the Inverted Hinge Loss (IHL) that aims to replace each token to unlearn with the next most-probable token, and show that IHL enables fast and stable tuning by resolving the issues of GA.
-Second, we find that the low-rank of LoRA imposes regularization too strong for LLM unlearning, and forcing the model to fully unlearn data points through extensive tuning leads to suboptimal cost vs.
+Beyond its efficiency, the low-rank structure in LoRA also serves as a strong regularizer (Biderman et al., 2024), which we hypothesize aids LLM unlearning by stabilizing optimization and mitigating catastrophic forgetting of retained knowledge.
+However, the empirical effects of LoRA in the context of LLM unlearning remain largely unexplored.
+In this paper, we present the first in-depth study of LLM unlearning under the low-rank adaptation paradigm and introduce two novel techniques for robust and parameter-efficient knowledge unlearning.
+First, we analyze the derivatives of GA to highlight its drawbacks such as (1) unnecessary forgetting induced by indiscriminately increasing the probability of all tokens and (2) unstable optimization due to the unboundedness of maximizing the next-token prediction loss.
+To tackle these challenges, we introduce the Inverted Hinge Loss (IHL), which replaces each unwanted token with the next most-probable one, and demonstrate that IHL facilitates stable unlearning by resolving the limitations of GA.
+Second, we find that the low-rank regularization in LoRA is too strong when unlearning with IHL, leading to suboptimal cost vs.
 post-unlearning performance trade-offs.
-To accelerate unlearning in a controlled manner, we propose Fisher-Initialization of Low-rank Adapters (FILA), which data-adaptively assigns parameters responsible for generating unwanted information to adapters prior to tuning by decomposing the pretrained parameters weighted by the relative Fisher-information matrix.
-Our extensive experiments on unlearning samples in the Training Data Extraction Challenge dataset with GPT-Neo models as well as removing knowledge of fictitious author profiles using the TOFU benchmark with Phi-1.5B and Llama2-7B models consistently demonstrate that our proposed IHL combined with FILA significantly outperforms previous baselines in both efficiency and post-unlearning downstream performance.
+In response, we propose Fisher-Initialization of Low-rank Adapters (FILA), which data-adaptively assigns parameters responsible for generating unwanted information to adapters prior to tuning by decomposing the pretrained parameters weighted by the relative Fisher-information matrix (Figure 1).
+Experiments on the Training Data Extraction Challenge dataset (GPT-Neo) and the TOFU benchmark (Phi-1.5B, Llama2-7B) demonstrate that IHL combined with FILA outperforms existing baselines in both efficiency and post-unlearning performance.
+In summary, our main contributions are as follows:
+  • We analyze the shortcomings of GA—unbounded optimization and unnecessary forgetting—through its derivative and propose IHL to address these issues.
+  • We introduce FILA, a method to accelerate unlearning by data-adaptively assigning parameters responsible for unwanted information to low-rank adapters before tuning.
+  • We demonstrate that IHL combined with FILA outperforms previous baselines in terms of both efficiency and post-unlearning performance (Figure 2).
 Figure 2: Compute cost for successful unlearning vs.
 post-unlearning downstream performances.
-We unlearn 32 randomly sampled sequences from the Training Data Extraction Challenge from GPT-Neo-125M.
+We unlearn 32 randomly sampled sequences in the Training Data Extraction Challenge from GPT-Neo-125M.
 Each point represents a different forget set and LoRA rank (if used).
@@
 Unlearning via gradient differences (GD) with vanilla LoRA leads to significant loss in performance compared to full-parameter GD unlearning due to lack of plasticity.
-However, our proposed method using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the cost-efficiency of LoRA.
+However, our proposed method using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the parameter-efficiency of LoRA.
 2 Related Work
 Machine Unlearning.
-The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pretrained model to discard information acquired from a specific subset of data previously used for pretraining.
-Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2023; Mehta et al., 2022; Chundawat et al., 2023; Cha et al., 2024).
+The primary objective of machine unlearning is to adapt a pretrained model to discard information acquired from a specific subset of data previously used during pretraining, with active research focused on image classification (Cao & Yang, 2015; Golatkar et al., 2020; Tarun et al., 2023; Mehta et al., 2022; Chundawat et al., 2023; Cha et al., 2024).
 Recently, its significance has grown notably with Large Language Models (LLMs) due to crucial need for managing unintended memorization of pretraining data intrinsic to LLMs (Si et al., 2023; Yao et al., 2024b).
-Several machine unlearning algorithms have been proposed for LLMs, with most falling under the category of parameter optimization-based methods (Si et al., 2023).
-First, Wang et al.
-(2023) introduced the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets.
-Chen & Yang (2023) proposed the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model intact.
-Liu et al.
-(2024) developed a two-stage framework: the harmful knowledge acquisition stage identifies and captures harmful knowledge within the model, and the knowledge negation stage removes that knowledge.
-While these approaches have demonstrated impressive unlearning results across various datasets, they are limited in applicability either by the need to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) or by the need to infer from a secondary model for knowledge distillation (Wang et al., 2023; Liu et al., 2024).
-Another category of LLM unlearning methods are tuning-free model-editing methods, an example of which leverages task arithmetic to suppress harmful text (Ilharco et al., 2023; Wu et al., 2023).
-While free from cost of tuning, these methods fails to show sufficient level of unlearning (Ilharco et al., 2023) or incur higher computational costs for detecting privacy neurons (Wu et al., 2023).
-In contrast, inspired by image classifier unlearning literature (Golatkar et al., 2020; Cha et al., 2024), Jang et al.
-(2023) adopted Gradient Ascent (GA) to unlearn from LLMs by maximizing the next-token prediction loss for all sequences in the forget set.
-Unlike in image classification, they demonstrate that LLMs show negligible forgetting of retained knowledge when using GA on the forget data.
-Consequently, they show that effective unlearning for LLMs is achievable solely using the forget dataset, while preserving general LLM knowledge and maintaining performance on downstream tasks.
-More recently, Yao et al.
-(2024a) demonstrated that incorporating GA with gradient descent on in-distribution data further enhances robustness against hyperparameters.
+Several machine unlearning algorithms for LLMs focus on parameter optimization (Si et al., 2023).
+For example, Wang et al.
+(2023) introduced Knowledge Gap Alignment, using knowledge distillation between models trained on different datasets.
+Chen & Yang (2023) proposed an unlearning layer to selectively remove specific knowledge while preserving other parameters, while Liu et al.
+(2024) developed a two-stage framework to capture and negate harmful knowledge.
+However, these approaches are limited by the need to retain large datasets (Wang et al., 2023; Chen & Yang, 2023) or rely on secondary models for distillation (Wang et al., 2023; Liu et al., 2024).
+Model-editing methods, such as task arithmetic for suppressing harmful content (Ilharco et al., 2023; Wu et al., 2023), avoid substantial costs but show limited effectiveness in unlearning.
+In contrast, Jang et al.
+(2023) used Gradient Ascent (GA) for LLM unlearning by maximizing the next-token prediction loss on the forget data, effectively unlearning while preserving model performance.
+Following this, GA has become a standard baseline, with Yao et al.
+(2024a) improving its robustness by combining GA with gradient descent on in-distribution data.
 Parameter-Efficient Fine-Tuning.
-Full-finetuning LLMs to adapt towards specific downstream tasks and instructions incurs an intractable computational cost due to the large model size and complexity.
-To alleviate this burden, previous work has developed Parameter-Efficient Fine-Tuning (PEFT) methods that adapt only a small portion of LLM parameters while freezing the other pretrained parameters intact (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023).
-Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2022; Zhang et al., 2023; Yeh et al., 2023; Kopiczko et al., 2024; yang Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter.
-A key benefit of LoRA is that the adapter weights can be merged seamlessly to pretrained parameters after fine-tuning, such that the post-adaptation LLM shares the same inference cost as the base pretrained LLM.
-While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposed to initialize LoRA weights using the principal singular vectors and values of the linear weights.
+Fine-tuning LLMs for specific tasks is computationally expensive due to their large parameter sizes.
+To address this, Parameter-Efficient Fine-Tuning (PEFT) methods adapt only a small subset of parameters while keeping the pretrained ones frozen (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023).
+Inspired by the small intrinsic rank of LLMs (Li et al., 2018; Aghajanyan et al., 2021), LoRA and its derivatives add low-rank adapters to the model’s linear layers (Hu et al., 2022; Zhang et al., 2023; Yeh et al., 2023; Kopiczko et al., 2024; yang Liu et al., 2024).
+These adapters can be merged with pretrained parameters after fine-tuning, maintaining the original inference cost.
+While most methods use random initialization for LoRA adapters, PiSSA (Meng et al., 2024) suggests initializing them using the principal singular vectors and values of the linear weights.
+Although various methods in machine unlearning and parameter-efficient fine-tuning for LLMs have been discussed, this paper focuses on analyzing the inherent issues of GA, introducing a novel unlearning loss function to overcome these issues, and exploring parameter-efficient unlearning methods that do not require full fine-tuning.
+These areas have not been adequately addressed in previous studies, highlighting the contributions of our work.
 3 Proposed Method
-In this section, we introduce useful notation and provide background information on unlearning with Gradient Ascent (GA) and Low-Rank Adaptation (LoRA).
-We then analyze derivatives of GA and highlight its inherent issues.
-Building upon this analysis, we present the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA.
-Lastly, we illustrate the method of initializing LoRA data-adaptively via Fisher-information, thereby enabling cost-efficient unlearning.
 3.1 Preliminaries
@@
 2023).
-GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
-$\mathcal{L}_{\text{GA}}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t})).$ (1)
-In practice, the log-likelihood is computed using cross-entropy loss and thus we GA essentially minimizes the Negative Cross-Entropy (NCE) loss.
+GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss
+$\mathcal{L}_{\text{GA}}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$ (1)
+In practice, the log-likelihood is computed using cross-entropy loss and thus GA essentially minimizes the Negative Cross-Entropy (NCE) loss.
 Therefore, GA maximizing the next-token prediction loss involves unbounded optimization, leading to an ill-posed process with unstable tuning.
@@
 Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank (Aghajanyan et al., 2021), LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter.
-In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes:
-$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}.$
+In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes
+$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}$
 During fine-tuning, the original weight ${\bm{W}}$ is kept frozen and only the low-rank factors ${\bm{A}}$ and ${\bm{B}}$ are updated via gradient descent.
@@
 $\frac{\partial\log\left(p_{\theta}(x_{t}|x_{<t})\right)}{\partial y_{t}^{(v)}}=\begin{cases}1-p_{\theta}(x_{t}|x_{<t})&\text{if }v=x_{t}\\ -p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
-From this derivative of GA, we can interpret its unlearning mechanism: GA decreases the prediction score corresponding to the true token $x_{t}$, while increasing the prediction score for all other tokens $v\neq x_{t}$ in the vocabulary.
-Consequently, we hypothesize that GA suffers from the following problems: 1) the objective is unbounded with the possibility of divergence during optimization, 2) gradient updates are inefficient due to gradients spreading across all other logits with large vocabulary sizes, and 3) there is unnecessary degradation of generative performance due to increasing the logits for all other tokens regardless of textual coherence.
+From this derivative of GA, we can interpret its unlearning mechanism: given the prefix $x_{<t}$, GA reduces the prediction score of the true token $x_{t}$ in proportion to $1-p_{\theta}(x_{t}|x_{<t})$ while increasing the scores of other tokens (i.e., $v\neq x_{t}$) by $p_{\theta}(v|x_{<t})$.
+This process effectively shifts the model’s prediction for $x_{<t}$ away from the true token $x_{t}$, thereby achieving unlearning.
+However, we can confirm that GA suffers form the following problems during unlearning: (1) Gradient spread, where reducing the score of $x_{t}$ while increasing the scores of all other tokens leads to inefficient unlearning in large vocabularies by predominantly boosting other tokens; (2) Unbounded loss, where minimizing $\log(p_{\theta}(x_{t}|x_{<t}))$ through maximizing cross-entropy loss introduces a risk of divergence due to the unbounded nature of entropy; and (3) Degradation of generative performance, where GA applies uniform gradient updates (i.e.
+increasing the scores of other tokens) to all sequences in the forget set $\mathcal{D}_{f}$, despite each sequence requiring a unique number of updates for unlearning.
+This redundancy can cause degrade the model’s generative capabilities, resulting in catastrophic forgetting.
 Inverted Hinge Loss.
 To cope with aforementioned limitations of GA, we aim to design a new loss function that achieves effective unlearning by decreasing the prediction score of the true token, while focusing gradient updates on only a minimal number of viable replacements for the ground-truth token.
-Inspired by the Hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge loss (IHL) as below:
-$\mathcal{L}_{\text{IHL}}({\bm{x}})=\max\left(0,1+p_{\theta}(x_{t}|x_{<t})-\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))\right)$
+Inspired by the Hinge Loss (Cortes & Vapnik, 1995), we devise Inverted Hinge Loss (IHL) as
+$\mathcal{L}_{\text{IHL}}({\bm{x}})=1+p_{\theta}(x_{t}|x_{<t})-\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))$
 As the probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function, the derivative of $\mathcal{L}_{\text{IHL}}({\bm{x}})$ with respect to $y_{t}^{(v)}$ is:
-$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}}.\end{cases}$
-Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IHL}}({\bm{x}})\neq 0$.
-Note that $\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IHL}}({\bm{x}})=0$.
-The detailed derivation can be found in Appendix A.
-The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly illustrate how the IHL addresses the shortcomings of GA in knowledge unlearning for LLMs.
-First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with opposite sign).
-This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly but also prevents spreading out of gradients.
-During this process, the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ increase slowly in proportion to the difference between the predictions for $x_{t}$ and $v^{\star}$.
-Second, once knowledge unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$), the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease.
-This not only prevents unnecessary forgetting but also results in a bounded loss function, allowing stable unlearning via stochastic gradient descent.
+$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$
+where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$. The detailed derivation can be found in Appendix A.
+The above derivative clearly illustrates how the IHL addresses the shortcomings of GA in knowledge unlearning for LLMs.
+First, IHL mitigates gradient spread by ensuring that the gradients primarily focus on the true token $x_{t}$ and its competitive token $v^{\star}$, without excessively boosting irrelevant tokens in large vocabularies.
+For example, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})>p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with opposite sign) and exceeds that of other tokens (i.e., $v\neq x_{t}\ \text{and}\ v\neq v^{\star}$).
+This ensures efficient and targeted unlearning while avoiding the unnecessary spread of gradients across irrelevant tokens.
+Second, IHL resolves the issue of unbounded loss by defining a bounded loss function, ensuring that the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease once unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$).
+This bounded loss prevents instability and divergence during unlearning.
+Lastly, IHL prevents the degradation of generative performance by adapting gradient updates based on the prediction scores of $x_{t}$ and $v^{\star}$, alleviating redundant updates for other tokens (i.e., $v\neq x_{t}\ \text{and}\ v\neq v^{\star}$).
+Compared to GA, this approach reduces redundant updates for other tokens in each sequence of the forget set $\mathcal{D}_{f}$, thereby preserving the model’s generative capabilities while achieving effective and stable unlearning.
 3.4 FILA: a Novel LoRA Initialization for LLM Unlearning
 Motivation.
-Biderman et al.
-(2024) discovered that, while it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios.
-While this stability is beneficial to retaining knowledge of ${\mathcal{D}}_{r}$ during unlearning, it also imposes a strong burden under the strict objective to remove knowledge of ${\mathcal{D}}_{f}$ completely.
-As a result, vanilla LoRA unlearning requires large number of model updates through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in Figure 2).
-Drawing inspiration from PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are relatively more important to ${\mathcal{D}}_{f}$ than to ${\mathcal{D}}_{r}$ would reinforce the model’s plasticity to forgetting ${\mathcal{D}}_{f}$, thereby accelerating the rate of unlearning and minimizing catastrophic forgetting on ${\mathcal{D}}_{r}$.
-In order to quantify the importance of parameters to particular data points, we leverage Fisher information matrices.
+While IHL effectively stabilizes the unlearning process by reducing the likelihood of unwanted tokens in a controlled manner, we empirically find that combining IHL with LoRA naïvely leads to requiring large number of unlearning iterations to fully forget samples in ${\mathcal{D}}_{f}$.
+Naturally, an increasing number of updates with ${\mathcal{D}}_{f}$ comes with the risk of overfitting, which for unlearning, may result in significant loss of knowledge on ${\mathcal{D}}_{r}$.
+We hypothesize that this is due to the random initialization of LoRA weights together with its low-rank structure imposing too strong a regularization to handle precise gradients needed for proper minimization of IHL (Biderman et al., 2024).
+Drawing inspiration from PiSSA (Meng et al., 2024), we therefore aim to accelerate optimization of IHL under LoRA by extracting weights relatively more important to ${\mathcal{D}}_{f}$ than to ${\mathcal{D}}_{r}$ from each pretrained weight, then using them to initialize LoRA weights prior to unlearning.
+We conjecture that this approach reinforces the model’s plasticity to forget ${\mathcal{D}}_{f}$ as well as its stability on keeping knowledge on ${\mathcal{D}}_{r}$.
+The remainder of this section presents how parameter importances are measured via Fisher information, and how low-rank adapter weights are initialized based on the measured importances.
 Parameter Importances via Fisher Information.
-The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information a dataset ${\mathcal{D}}$ provides on model parameters $\theta$.
-More concretely, ${\bm{F}}_{\theta}$ is computed as the second cross-moments of first partial derivatives of the log-likelihood of ${\mathcal{D}}$ (left of Eq.
-2).
-However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead.
+The Fisher information matrix ${\bm{F}}_{\theta}({\mathcal{D}})$ is defined as the variance of the partial derivative of the log-likelihood of data ${\mathcal{D}}$ with respect to the model parmeter $\theta$ (left of Equation 2).
+Intuitively, the matrix can be considered a measurement on how much the model output changes following a small change on its parameter weight.
+However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}({\mathcal{D}})$ instead.
 In the context of LLMs, this can be computed as:
@@
 $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$).
-Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to identify parameters that are important exclusively for ${\mathcal{D}}_{f}$ and not for ${\mathcal{D}}_{r}$.
+Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times k}$ as an importance metric to identify parameters that are important exclusively for ${\mathcal{D}}_{f}$ and not for ${\mathcal{D}}_{r}$.
 While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus use $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$ rather than $\hat{{\bm{F}}}_{\bm{W}}^{f}$.
@@
 with $\mathds{1}\in\mathbb{R}^{k}$ and $\texttt{diag}(\cdot)$ indicating the all-one vector and the vector diagonalization function, respectively.
-Unlike general WLRA, this row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
+Unlike general WLRA, this row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\texttt{diag}((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{\frac{1}{2}}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{-\frac{1}{2}}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
 Given this solution, we use ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights.
 To ensure that the model behavior remains the same after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}^{*}{\bm{A}}^{*}$.
-Intuitively, our Fisher-weighted Initialization of Low-rank Adapters (FILA) extracts parameters that are important for generating ${\mathcal{D}}_{f}$, but not for generating ${\mathcal{D}}_{r}$, such that LoRA tuning can be focused on erasing knowledge relevant to ${\mathcal{D}}_{f}$ while keeping information regarding ${\mathcal{D}}_{r}$.
+Intuitively, our Fisher-weighted Initialization of Low-rank Adapters (FILA) extracts parameters that are important for generating ${\mathcal{D}}_{f}$ but not for generating ${\mathcal{D}}_{r}$, such that LoRA tuning can be focused on erasing knowledge relevant to ${\mathcal{D}}_{f}$ while keeping information regarding ${\mathcal{D}}_{r}$.
 3.5 Final Loss Function for LLM Unlearning
-In summary, we perform unlearning on the model $\bm{\Theta}=\bm{\theta}\cup\bm{\theta}_{\text{FILA}}$, consisted of original pretrained weights $\bm{\theta}$ and the FILA-initialized low-rank adapter weights for each linear layer $\bm{\theta}_{\text{FILA}}=\{{\bm{A}}^{*}_{\ell},{\bm{B}}^{*}_{\ell}\}_{\ell=1}^{L}$, where $L$ represents the number of layers tuned via LoRA.
+In summary, we perform unlearning on the model $\bm{\Theta}=\bm{\theta}\cup\bm{\theta}_{\text{FILA}}$ consisted of original pretrained weights $\bm{\theta}$ and FILA-initialized low-rank adapter weights for each linear layer $\bm{\theta}_{\text{FILA}}=\{{\bm{A}}^{*}_{\ell},{\bm{B}}^{*}_{\ell}\}_{\ell=1}^{L}$, where $L$ represents the number of layers tuned via LoRA.
 Additionally, we incorporate GD, which utilizes the auxiliary retain set $\mathcal{D}_{r}$.
@@
 $\operatorname*{minimize}_{\theta_{\text{FILA}}}\sum_{{\bm{x}}_{r}\in{\mathcal{D}}_{f},{\bm{x}}_{f}\in{\mathcal{D}}_{r}}\mathcal{L}_{\text{IHL}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (3)
-In practice, training (unlearning) for the LLM model is conducted by minimizing Eq.3 through stochastic gradient descent.
+In practice, training (unlearning) for the LLM model is conducted by minimizing Equation 3 through stochastic gradient descent.
 4 Experiments
@@
 For brevity, we present results from additional experiments such as continual unlearning in Appendix D.
-Table 1: Evaluation results on LM-Evaluation, Dialogue datasets before and after unlearning samples from the TDEC dataset.
-For all runs under LoRA, we use rank 16 and GD is used as the baseline.
-The "+IHL" results refer to experiments where NCE loss within GD is replaced with our proposed IHL (i.e., Equation 3), and "+FILA" indicates results where FILA is used to initialize LoRA.
+Table 1: Evaluation results on reasoning and generative capabilities before and after unlearning samples from the TDEC dataset.
+For each model, we test three unlearning objectives (i.e.
+GA, GD, IHL) with full-parameter tuning, and four methods using LoRA with rank 16.
+IHL denotes replacing the GA loss term in GD with our IHL, and +FILA denotes initializing LoRA weights with FILA.
+Changes in performance after unlearning are in parentheses, with best results under LoRA-based unlearning in bold.
 Model Method Params. Epochs EL10 MA Reasoning Dialogue Pile
 (%) (%$\downarrow$ (%$\downarrow$ (Acc)$\uparrow$ (F1$\uparrow$ (PPL)$\downarrow$
-Before - - 30.9 77.4 43.4 9.4 17.8
-GA 100.0 17.2 1.0 27.4 39.9 2.6 577.8
-GPT-Neo GD 4.6 0.7 24.9 42.4 5.9 54.2
-125M LoRA 1.6 8.6 0.3 20.6 40.8 2.5 129.4
-+ IHL 11.4 0.4 22.7 41.9 6.0 32.9
-     + FILA 6.0 0.3 23.9 42.2 10.1 24.0
-Before - - 67.6 92.2 49.8 11.5 11.5
-GA 100.0 13.8 1.9 30.4 49.7 8.5 15.8
-GPT-Neo GD 12.8 2.2 30.9 48.4 12.7 10.8
-1.3B LoRA 0.8 19.3 1.7 31.4 45.0 9.7 31.8
-+ IHL 20.0 1.7 44.6 47.1 10.2 14.9
-     + FILA 13.0 0.5 29.6 48.3 12.1 14.7
-Before - - 70.4 93.4 52.3 11.5 10.4
-GA 100.0 10.8 1.6 31.0 51.9 11.1 17.9
-GPT-Neo GD 8.0 0.7 28.3 44.0 12.7 17.9
-2.7B LoRA 0.7 14.0 0.1 20.4 45.9 6.7 61.1
-+ IHL 17.8 0.0 26.7 49.6 8.5 22.2
-     + FILA 10.3 0.1 28.5 49.6 10.7 16.0
+GPT-Neo 125M Before - - 30.9 77.4 43.4 9.4 17.8
+GA 100.0 17.2 1.0 27.4 39.9 (-3.5) 2.6 (-6.8) 577.8 (+560.0)
+GD 4.6 0.7 24.9 42.4 (-1.0) 5.9 (-3.5) 54.2 (+36.4)
+IHL 17.2 0.7 29.2 42.3 (-1.1) 10.3 (+0.9) 18.1 (+0.3)
+GD 1.6 8.6 0.3 20.6 40.8 (-2.6) 2.5 (-6.9) 129.4 (+111.6)
+IHL 11.4 0.4 22.7 41.9 (-1.5) 6.0 (-3.4) 32.9 (+15.1)
+GD+FILA 7.4 1.2 27.4 42.0 (-1.4) 6.5 (-2.9) 89.5 (+71.7)
+IHL+FILA 6.0 0.3 23.9 42.2 (-1.2) 10.1 (+0.7) 24.0 (+6.2)
+GPT-Neo 1.3B Before - - 67.6 92.2 49.8 11.5 11.5
+GA 100.0 13.8 1.9 30.4 49.7 (-0.1) 8.5 (-3.0) 15.8 (+4.3)
+GD 12.8 2.2 30.9 48.4 (-1.4) 12.7 (+1.2) 10.8 (-0.7)
+IHL 7.6 0.7 30.4 48.4 (-1.4) 12.5 (+1.0) 11.0 (-0.5)
+GD 0.8 19.3 1.7 31.4 45.0 (-4.8) 9.7 (-1.8) 31.8 (+20.3)
+IHL 20.0 1.7 44.6 47.1 (-2.7) 10.2 (-1.3) 14.9 (+3.4)
+GD+FILA 7.8 1.9 23.2 44.2 (-5.6) 5.5 (-6.0) 54.5 (+43.0)
+IHL+FILA 13.0 0.5 29.6 48.3 (-1.5) 12.1 (+0.6) 14.7 (+3.2)
+GPT-Neo 2.7B Before - - 70.4 93.4 52.3 11.5 10.4
+GA 100.0 10.8 1.6 31.0 51.9 (-0.4) 11.1 (-0.4) 17.9 (+7.5)
+GD 8.0 0.7 28.3 51.8 (-0.5) 12.7 (+1.2) 17.9 (+7.5)
+IHL 6.6 0.5 29.3 51.8 (-0.5) 12.9 (+1.4) 10.7 (+0.3)
+GD 0.7 14.0 0.1 20.4 45.9 (-6.4) 6.7 (-4.8) 61.1 (+50.7)
+IHL 17.8 0.0 26.7 49.6 (-2.7) 8.5 (-2.6) 22.2 (+11.8)
+GD+FILA 6.8 1.6 28.9 44.8 (-7.5) 9.3 (-2.2) 68.7 (+58.3)
+IHL+FILA 10.3 0.1 28.5 49.6 (-2.7) 10.7 (-0.8) 16.0 (+5.6)
 4.1 Training Data Extraction Challenge
 Experimental Setup.
-The Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) found to be easily extractable from a pretrained LLM.
+The Training Data Extraction Challenge (TDEC) dataset (Carlini et al., 2021) consists of 20k examples from the Pile dataset (Gao et al., 2020) found to be easily extractable from a pretrained LLM.
 For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$.
@@
 We exclude methods that require another auxiliary model (Wang et al., 2023; Liu et al., 2024) or the entire training data (Wang et al., 2023; Chen & Yang, 2023) from our baselines.
-Figure 3: Results from unlearning examples in the TDEC dataset from the GPT-Neo LLMs with three different sizes.
+Figure 3: Results from unlearning examples in the TDEC dataset on the GPT-Neo LLM family.
 Each row represents the performance averaged across datasets within each set of LLM capability tests: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better).
@@
 The dashed lines indicate the base model performance prior to unlearning.
-Unlearning with GD leads to significant loss in performance and also frequently fails to unlearn effectively even with large LoRA ranks.
-Replacing the negative cross-entropy (NCE) loss with our hinge loss (IHL) boosts retention of reasoning and generation capabilities, but still fails to unlearn in multiple cases.
-Using NCE loss with FILA initialization notably increases the rate of unlearning success, but at the cost of losing overall performance.
+GD leads to significant loss in performance and also fails to unlearn in some cases even with large LoRA ranks.
+Replacing the NCE loss in GD with IHL boosts retention of reasoning and generation capabilities, but still fails to unlearn in multiple cases.
+Running GD with FILA notably increases the rate of unlearning success, but at significant cost in overall performance.
 Using both IHL and FILA best minimizes post-unlearning performance degradation in all three aspects.
@@
 Mainly for 125M and 1.3B models, GA causes a significant decline in generative performance, whereas GD partially mitigates this decline and improves both Dialogue F1 scores and the Pile perplexity.
-Second, we find that simply applying GD with LoRA fails to achieve successful unlearning across all model sizes.
-Second, we find that imply running GD with LoRA significantly increases the number of epochs required for successful unlearning.
-While enjoying great parameter-efficiency by tuning only about 0.7% to 1.6% of the total parameters when using LoRA, its application to unlearning with GD results in significant loss in overall language capability, especially on generative tasks.
-Lastly, using our proposed IHL in conjunction with FILA achieves the forgetting criteria in fewest epochs while best preserving LLM performance.
-Although IHL with LoRA alone retains knowledge better than the baseline, it requires more epochs to meet the forgetting criteria.
-This is resolved when used together with FILA, and the number of required epochs is significantly reduced, reducing catastrophic forggetting of previously acquired knowledge even further.
-Figure 4: Results from unlearning examples from TDEC dataset using LoRA with rank 32 to adapt sets of layers on GPT-Neo 2.7B.
-The marker shapes and colors are used similarly as in Figure 3.
-Based on the rate of unlearning success, tuning FFN layers (e.g., FFN, QVFFN) is more receptive to targeted knowledge removal compared to tuning attention layers (e.g., QV, QKVO).
+Second, we find that simply applying GD with LoRA fails to unlearn effectively across all model sizes.
+While enjoying great parameter-efficiency by tuning only about 0.7% to 1.6% of the total parameters when using LoRA, its application to unlearning with GD results in large loss in overall language capability, especially on generative tasks.
+Third, replacing GA in GD with IHL leads to performance gains in both full-parameter and LoRA-based unlearning, but requires larger number of epochs for successful unlearning than GD especially when confined to low-rank weight changes.
+However, this is resolved when IHL is used together with FILA, which significantly reduces the number of required epochs.
+Note that using FILA with GD also reduces the number of epochs required, but worsens downstream performance in many cases, as FILA essentially accelerates the divergent behavior of GA in GD.
+In essence, whether the ability of FILA to accelerate tuning also translates to benefits in downstream performance depends on the loss function being optimized.
+On the other hand, when paired with our bounded IHL, FILA leads to better retention of LLM performance, showcasing the strong synergy of our approach that enjoys not only the stability of IHL but also the speedup from FILA.
 4.2 Analysis
@@
 Furthermore, integrating FILA with IHL achieves the best post-unlearning performance across all LoRA target module combinations.
+Figure 4: Results from unlearning examples from TDEC dataset using LoRA with rank 32 to adapt sets of layers on GPT-Neo-125M.
+The marker shapes and colors are used similarly as in Figure 3.
+Based on the rate of unlearning success, tuning FFN layers (e.g., FFN, QVFFN) is more receptive to targeted knowledge removal compared to tuning attention layers (e.g., QV, QKVO).
 Cost-efficiency of the proposed method.
@@
 A high p-value indicates high distributional similarity between the unlearned model and the reference model, thus implying strong forgetting.
-To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregated model performance.
-Further details on the dataset and evaluation pipeline can be found in  Maini et al.
-(2024).
+To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregation of the probability of correct answers, ROUGE-L scores, and Truth Ratio of correct answers vs.
+incorrect answers on questions from three datasets pertaining to the retain set of fictitious authors, real authors, and world facts.
 Following the original paper of TOFU, we prepare two base models by finetuning Phi-1.5B and Llama2-7B on TOFU for 5 epochs with learning rates 2e-5 and 1e-5, respectively.
-We then unlearn using two baselines (GA and GD) and our two methods (IHL and IHL+FILA) using LoRA adapters of rank 4, 8, 16, or 32.
+We then unlearn with various methods using LoRA adapters of rank 4, 8, 16, or 32.
+While we mainly compare our methods against GD, we also compare IHL and FILA against existing unlearning methods such as KL (Maini et al., 2024), DPO (Rafailov et al., 2024), and NPO (Zhang et al., 2024), results from which can be found in Appendix D.
 For unlearning, we use a learning rate of 2e-4 if our base model is from Phi-1.5B and 1e-4 for Llama2-7B.
@@
 (b) Llama2-7B
-[…]
+Figure 5: TOFU results using Phi-1.5B and Llama2-7B models.
+Each row corresponds to unlearning a different forget set (1%, 5%, or 10%), and each column uses a distinct LoRA rank between 4 and 32.
+The relative size of markers represent the number of epochs.
+Ideally, the unlearning curves should start from the pretrained model ($\blacksquare$) and approach towards the reference model tuned on the retain set only ($\bigstar$) as unlearning progresses.
+Both GD and GD+FILA suffers from significant loss of model utility due to using GA for unlearning.
+Replacing GA with IHL largely retains model utility, then initializing LoRA adapters with FILA significantly boosts the unlearning efficiency of IHL.
 Results.
@@
 With regards to our proposed method, we find that replacing the NCE loss in GD with our IHL better retains model utility across all LoRA ranks and forget set sizes, as curves are more aligned straight-up towards the reference point with negligible shift in model utility.
-This stability comes at the cost of unlearning efficiency, however, as the rate at which the LLM forgets ${\mathcal{D}}_{f}$ is slower with IHL due to IHL decreasing the likelihood of the unwanted token by increasing the likelihood of only one other most-possible token each time in a controlled manner.
-Nonetheless, initializing LoRA adapters with FILA largely alleviates this issue and enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.
+This stability comes at the cost of unlearning efficiency, however, as randomly initialized LoRA weights are unable to effectively represent weight changes required to decrease IHL.
+Nonetheless, initializing LoRA adapters with FILA largely alleviates this issue and significantly enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.
 Interestingly, we find the prior weight assignment via FILA can lead to excessive unlearning in some cases (e.g., unlearning 10% forget set with ranks 8 or 16 on Llama2-7B), with model updates reducing the forget quality after reaching the upper bound at zero.
@@
 $\frac{\partial p_{\theta}(x_{t}|x_{<t})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))&\text{if }v=x_{t}\\ -p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
-Based on the derivative of $p_{\theta}(x_{t}|x_{<t})$ above , we can calculate the derivative of $\mathcal{L}_{\text{IH}}$.
+Based on the derivative of $p_{\theta}(x_{t}|x_{<t})$ above , we can calculate the derivative of $\mathcal{L}_{\text{IHL}}$.
 Firstly, for convenience, we define $p_{t}=p_{\theta}(x_{t}|x_{<t})$ and $\hat{p}_{t}=\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))$.
 The loss function can be rewritten as:
-$\mathcal{L}_{\text{IH}}({\bm{x}})=\max\left(0,1+p_{t}-\hat{p}_{t}\right)$
-To calculate the derivative of $\mathcal{L}_{\text{IH}}$, we need to consider three cases: 1) when $v=x_{t}$, 2) when $v=v^{\star}$ where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$, 3) when $v\neq x_{t}$ and $v\neq v^{\star}$.
-In the case where $01+p_{t}-\hat{p}_{t}>$ 0, using the derivative of $p_{\theta}(x_{t}|x_{<t})$ mentioned earlier, the derivative of $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ is as follows:
-$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(x_{t})}}$ $=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
+$\mathcal{L}_{\text{IHL}}({\bm{x}})=1+p_{t}-\hat{p}_{t}$
+To calculate the derivative of $\mathcal{L}_{\text{IHL}}$, we need to consider three cases: 1) when $v=x_{t}$, 2) when $v=v^{\star}$ where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$, 3) when $v\neq x_{t}$ and $v\neq v^{\star}$.
+Using the derivative of $p_{\theta}(x_{t}|x_{<t})$ mentioned earlier, the derivative of $\mathcal{L}_{\text{IHL}}$ with respect to $y_{t}^{(v)}$ is as follows:
+$\frac{\partial\mathcal{L}_{\text{IHL}}}{\partial y_{t}^{(x_{t})}}$ $=\frac{\partial}{\partial y_{t}^{(x_{t})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
 $=p_{\theta}(x_{t}|x_{<t})(1-p_{\theta}(x_{t}|x_{<t}))+p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v^{\star}|x_{<t})$
 $=p_{\theta}(x_{t}|x_{<t})\left(1-p_{\theta}(x_{t}|x_{<t})+p_{\theta}(v^{\star}|x_{<t})\right)$
-$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v^{\star})}}$ $=\frac{\partial}{\partial y_{t}^{(v^{\star})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
+$\frac{\partial\mathcal{L}_{\text{IHL}}}{\partial y_{t}^{(v^{\star})}}$ $=\frac{\partial}{\partial y_{t}^{(v^{\star})}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
 $=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})(1-p_{\theta}(v^{\star}|x_{<t}))$
 $=-p_{\theta}(v^{\star}|x_{<t})\left(1-p_{\theta}(v^{\star}|x_{<t})+p_{\theta}(x_{t}|x_{<t})\right)$
-$\frac{\partial\mathcal{L}_{\text{IH}}}{\partial y_{t}^{(v)}}$ $=\frac{\partial}{\partial y_{t}^{(v)}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
+$\frac{\partial\mathcal{L}_{\text{IHL}}}{\partial y_{t}^{(v)}}$ $=\frac{\partial}{\partial y_{t}^{(v)}}\left(1+p_{\theta}(x_{t}|x_{<t})-p_{\theta}(v^{\star}|x_{<t})\right)$
 $=-p_{\theta}(x_{t}|x_{<t})\cdot p_{\theta}(v|x_{<t})+p_{\theta}(v^{\star}|x_{<t})\cdot p_{\theta}(v|x_{<t})$
 $=p_{\theta}(v|x_{<t})\left(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})\right)$
-Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $1+p_{t}-\hat{p}_{t}\leq 0$.
-In summary, the derivatives of the loss function $\mathcal{L}_{\text{IH}}$ with respect to $y_{t}^{(v)}$ for the three cases are:
-$\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$
+In summary, the derivatives of the loss function $\mathcal{L}_{\text{IHL}}$ with respect to $y_{t}^{(v)}$ for the three cases are:
+$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$
 Appendix B Evaluation Metrics
@@
 Task 0 refers to the result before unlearning.
+D.2 TOFU Results with Additional Baselines
+We compare our IHL and FILA methods against three additional existing unlearning baselines: KL (Maini et al., 2024) uses GA to forget samples in ${\mathcal{D}}_{f}$ while minimizing the Kullback-Leibler (KL) divergence between representations of retain samples in ${\mathcal{D}}_{r}$ output by the unlearned model and those from the pretrained base model.
+Instead of GA, Direct Preference Optimization (DPO; Rafailov et al.
+(2024)) performs unlearning by training the model to output variants of “I don’t know” when given a sample in ${\mathcal{D}}_{f}$.
+NPO (Zhang et al., 2024) is an approach similar to GA, but with adaptive weighting on the gradients such that it alleviates the divergent behavior of GA.
+Note that both DPO and NPO are regularized by the LM-loss on ${\mathcal{D}}_{r}$.
+We use the same hyperparameterization (e.g., learning rate and effective batch size) as in our main results.
+Figure 7 shows the results.
+We find that all three baselines lead to significant decrease in model utility, while IHL+FILA shows negligible change.
+(a) Phi-1.5B
+(b) Llama2-7B
+Figure 7: TOFU results using Phi-1.5B and Llama2-7B models.
+Each row corresponds to unlearning a different forget set (1%, 5%, or 10%), and each column uses a distinct LoRA rank between 4 and 32.
+The relative size of markers represent the number of epochs.
+Ideally, the unlearning curves should start from the pretrained model ($\blacksquare$) and approach towards the reference model tuned on the retain set only ($\bigstar$) as unlearning progresses.
+Our method IHL+FILA outperforms existing KL- and preference optimization-based unlearning approaches in both model utility and forget quality.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.93 0.93 0.48 1.15e-17 0.92 0.92 0.48 0.41 0.37 0.45 0.75 0.41 0.50 0.52
+Retain90 0.43 0.14 0.63 1.00e+00 0.91 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.49 0.52
+TOFU Forget01
+KL 0.96 0.92 0.48 7.37e-05 0.92 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.50 0.52
+DPO 0.96 0.93 0.48 8.87e-05 0.92 0.92 0.48 0.44 0.38 0.45 0.75 0.41 0.50 0.52
+NPO 0.96 0.92 0.48 6.11e-05 0.92 0.92 0.48 0.43 0.37 0.45 0.76 0.41 0.50 0.52
+GA 0.96 0.92 0.48 6.11e-05 0.92 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.50 0.52
+GA+FILA 0.04 0.00 0.76 1.07e-03 0.06 0.00 0.21 0.01 0.27 0.29 0.02 0.29 0.30 0.00
+GD 0.96 0.93 0.48 7.37e-05 0.92 0.92 0.48 0.42 0.38 0.45 0.76 0.41 0.50 0.52
+IHL 0.96 0.93 0.48 4.17e-05 0.92 0.92 0.48 0.43 0.38 0.45 0.75 0.41 0.50 0.52
+GD+FILA 0.03 0.00 0.69 3.24e-02 0.06 0.00 0.20 0.00 0.26 0.31 0.03 0.27 0.31 0.00
+IHL+FILA 0.50 0.52 0.49 1.28e-04 0.83 0.89 0.49 0.37 0.38 0.45 0.73 0.41 0.50 0.51
+TOFU Forget05
+KL 0.62 0.49 0.51 2.90e-13 0.65 0.51 0.46 0.48 0.36 0.43 0.80 0.39 0.47 0.48
+DPO 0.43 0.82 0.51 2.17e-13 0.55 0.82 0.45 0.34 0.35 0.42 0.72 0.41 0.50 0.47
+NPO 0.62 0.48 0.51 4.87e-12 0.64 0.51 0.45 0.50 0.36 0.43 0.80 0.39 0.47 0.48
+GA 0.61 0.46 0.51 1.10e-11 0.63 0.48 0.45 0.46 0.36 0.43 0.80 0.39 0.46 0.47
+GA+FILA 0.09 0.00 0.73 3.69e-08 0.10 0.00 0.20 0.00 0.25 0.28 0.03 0.24 0.25 0.00
+GD 0.70 0.80 0.47 4.33e-15 0.79 0.88 0.48 0.37 0.38 0.45 0.72 0.41 0.50 0.50
+IHL 0.71 0.82 0.48 6.68e-14 0.83 0.90 0.48 0.37 0.38 0.45 0.73 0.41 0.49 0.50
+GD+FILA 0.12 0.00 0.72 4.33e-05 0.13 0.00 0.18 0.01 0.28 0.36 0.02 0.29 0.32 0.00
+IHL+FILA 0.45 0.37 0.50 1.44e-11 0.79 0.86 0.48 0.43 0.37 0.46 0.75 0.41 0.50 0.51
+TOFU Forget10
+KL 0.01 0.00 0.77 7.38e-15 0.01 0.00 0.16 0.00 0.26 0.24 0.00 0.27 0.25 0.00
+DPO 0.41 0.87 0.49 5.10e-17 0.67 0.88 0.47 0.33 0.36 0.43 0.73 0.41 0.49 0.48
+NPO 0.45 0.21 0.61 2.56e-05 0.45 0.20 0.38 0.35 0.34 0.39 0.71 0.37 0.43 0.37
+GA 0.01 0.00 0.76 2.06e-13 0.01 0.00 0.15 0.00 0.26 0.24 0.00 0.26 0.24 0.00
+GA+FILA 0.00 0.00 0.35 5.10e-17 0.00 0.00 0.25 0.00 0.25 0.38 0.00 0.23 0.32 0.00
+GD 0.37 0.19 0.53 2.55e-09 0.41 0.29 0.44 0.19 0.36 0.44 0.60 0.38 0.46 0.36
+IHL 0.53 0.67 0.49 2.43e-17 0.76 0.86 0.49 0.39 0.38 0.45 0.71 0.41 0.50 0.51
+GD+FILA 0.12 0.00 0.65 2.17e-06 0.11 0.00 0.23 0.00 0.30 0.30 0.03 0.26 0.28 0.00
+IHL+FILA 0.26 0.19 0.49 1.39e-12 0.75 0.85 0.50 0.36 0.39 0.49 0.67 0.41 0.51 0.51
+Table 2: Detailed TOFU Results using the Phi-1.5B model with LoRA rank = 4.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.93 0.93 0.48 1.15e-17 0.92 0.92 0.48 0.41 0.37 0.45 0.75 0.41 0.50 0.52
+Retain90 0.43 0.14 0.63 1.00e+00 0.91 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.49 0.52
+TOFU Forget01
+KL 0.91 0.91 0.48 6.11e-05 0.92 0.92 0.48 0.43 0.37 0.45 0.77 0.41 0.50 0.52
+DPO 0.96 0.93 0.49 8.87e-05 0.92 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.50 0.52
+NPO 0.92 0.91 0.48 6.11e-05 0.91 0.92 0.48 0.43 0.37 0.45 0.76 0.41 0.50 0.52
+GA 0.92 0.91 0.48 4.17e-05 0.92 0.92 0.48 0.43 0.37 0.45 0.77 0.41 0.50 0.52
+GA+FILA 0.00 0.00 0.65 2.72e-02 0.01 0.00 0.22 0.00 0.23 0.32 0.00 0.29 0.34 0.00
+GD 0.93 0.92 0.48 7.37e-05 0.92 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.50 0.52
+IHL 0.94 0.92 0.48 7.37e-05 0.92 0.92 0.48 0.43 0.38 0.45 0.75 0.41 0.50 0.52
+GD+FILA 0.01 0.00 0.65 4.55e-02 0.01 0.00 0.24 0.00 0.23 0.32 0.02 0.29 0.36 0.00
+IHL+FILA 0.47 0.33 0.51 3.37e-04 0.80 0.86 0.49 0.34 0.38 0.46 0.73 0.42 0.51 0.50
+TOFU Forget05
+KL 0.46 0.14 0.55 3.26e-10 0.45 0.15 0.44 0.38 0.34 0.42 0.71 0.36 0.43 0.35
+DPO 0.44 0.84 0.50 1.21e-13 0.64 0.85 0.46 0.34 0.36 0.42 0.72 0.41 0.50 0.48
+NPO 0.51 0.24 0.54 9.13e-11 0.50 0.25 0.43 0.48 0.34 0.42 0.78 0.37 0.44 0.41
+GA 0.45 0.12 0.55 3.26e-10 0.45 0.13 0.43 0.33 0.34 0.42 0.69 0.36 0.43 0.33
+GA+FILA 0.02 0.00 0.54 1.12e-09 0.02 0.00 0.22 0.00 0.23 0.33 0.00 0.25 0.30 0.00
+GD 0.47 0.49 0.49 3.70e-12 0.53 0.66 0.47 0.26 0.37 0.44 0.69 0.40 0.48 0.44
+IHL 0.62 0.74 0.48 1.21e-13 0.81 0.89 0.48 0.37 0.37 0.45 0.72 0.41 0.49 0.50
+GD+FILA 0.02 0.00 0.74 3.16e-07 0.02 0.00 0.16 0.00 0.26 0.32 0.02 0.29 0.33 0.00
+IHL+FILA 0.41 0.25 0.52 3.73e-09 0.77 0.86 0.48 0.38 0.37 0.47 0.71 0.41 0.50 0.50
+TOFU Forget10
+KL 0.01 0.00 0.76 1.06e-16 0.01 0.00 0.13 0.00 0.27 0.34 0.00 0.26 0.27 0.00
+DPO 0.39 0.87 0.49 1.15e-17 0.72 0.89 0.47 0.28 0.37 0.43 0.73 0.41 0.49 0.47
+NPO 0.46 0.24 0.63 1.66e-04 0.46 0.24 0.36 0.27 0.33 0.38 0.71 0.37 0.43 0.36
+GA 0.01 0.00 0.76 5.10e-17 0.01 0.00 0.13 0.00 0.27 0.34 0.00 0.26 0.27 0.00
+GA+FILA 0.00 0.00 0.39 2.43e-19 0.00 0.00 0.29 0.00 0.24 0.36 0.00 0.26 0.34 0.00
+GD 0.23 0.04 0.55 5.07e-06 0.27 0.13 0.43 0.07 0.35 0.45 0.44 0.38 0.45 0.22
+IHL 0.46 0.55 0.49 2.43e-17 0.77 0.85 0.49 0.38 0.38 0.45 0.69 0.41 0.50 0.50
+GD+FILA 0.10 0.00 0.60 1.66e-04 0.10 0.00 0.19 0.00 0.29 0.41 0.04 0.27 0.32 0.00
+IHL+FILA 0.18 0.09 0.52 1.40e-06 0.73 0.84 0.49 0.30 0.41 0.51 0.66 0.43 0.52 0.50
+Table 3: Detailed TOFU Results using the Phi-1.5B model with LoRA rank = 8.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.93 0.93 0.48 1.15e-17 0.92 0.92 0.48 0.41 0.37 0.45 0.75 0.41 0.50 0.52
+Retain90 0.43 0.14 0.63 1.00e+00 0.91 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.49 0.52
+TOFU Forget01
+KL 0.84 0.86 0.48 2.83e-05 0.91 0.92 0.48 0.46 0.37 0.45 0.75 0.41 0.49 0.53
+DPO 0.96 0.93 0.49 7.37e-05 0.91 0.92 0.48 0.42 0.38 0.45 0.76 0.42 0.51 0.52
+NPO 0.82 0.86 0.48 4.17e-05 0.91 0.92 0.48 0.44 0.37 0.45 0.76 0.41 0.49 0.52
+GA 0.84 0.86 0.48 6.11e-05 0.91 0.92 0.48 0.44 0.37 0.45 0.75 0.41 0.49 0.52
+GA+FILA 0.03 0.00 0.64 1.14e-02 0.01 0.00 0.19 0.01 0.22 0.27 0.01 0.29 0.29 0.00
+GD 0.88 0.88 0.48 7.37e-05 0.92 0.92 0.49 0.40 0.38 0.45 0.76 0.41 0.50 0.52
+IHL 0.88 0.89 0.48 1.28e-04 0.91 0.92 0.49 0.42 0.38 0.45 0.76 0.41 0.50 0.52
+GD+FILA 0.03 0.00 0.60 2.44e-01 0.02 0.00 0.19 0.00 0.22 0.27 0.01 0.30 0.36 0.00
+IHL+FILA 0.44 0.19 0.55 1.70e-03 0.75 0.84 0.49 0.37 0.39 0.47 0.72 0.43 0.53 0.51
+TOFU Forget05
+KL 0.21 0.00 0.69 1.53e-03 0.22 0.00 0.29 0.02 0.27 0.34 0.04 0.27 0.29 0.00
+DPO 0.43 0.85 0.50 6.68e-14 0.73 0.88 0.47 0.37 0.36 0.43 0.74 0.41 0.50 0.49
+NPO 0.46 0.16 0.58 1.10e-07 0.45 0.17 0.41 0.36 0.34 0.41 0.66 0.36 0.43 0.35
+GA 0.21 0.00 0.71 4.33e-05 0.21 0.00 0.26 0.01 0.26 0.32 0.04 0.26 0.27 0.00
+GA+FILA 0.02 0.00 0.46 5.96e-09 0.02 0.00 0.29 0.00 0.25 0.34 0.00 0.27 0.38 0.00
+GD 0.40 0.19 0.52 3.73e-09 0.43 0.39 0.45 0.12 0.35 0.41 0.53 0.37 0.45 0.32
+IHL 0.52 0.59 0.49 2.12e-12 0.79 0.87 0.48 0.38 0.38 0.45 0.71 0.41 0.49 0.50
+GD+FILA 0.08 0.00 0.69 1.81e-05 0.07 0.00 0.17 0.01 0.25 0.33 0.05 0.27 0.27 0.00
+IHL+FILA 0.36 0.11 0.57 5.03e-06 0.75 0.84 0.49 0.43 0.37 0.47 0.71 0.42 0.51 0.52
+TOFU Forget10
+KL 0.01 0.00 0.70 1.07e-13 0.01 0.00 0.14 0.00 0.29 0.41 0.00 0.25 0.35 0.00
+DPO 0.32 0.86 0.48 5.40e-18 0.76 0.90 0.48 0.32 0.37 0.43 0.72 0.41 0.49 0.48
+NPO 0.45 0.25 0.65 4.69e-04 0.45 0.27 0.35 0.30 0.33 0.37 0.69 0.37 0.42 0.37
+GA 0.01 0.00 0.71 1.46e-14 0.01 0.00 0.14 0.00 0.29 0.41 0.00 0.25 0.35 0.00
+GA+FILA 0.00 0.00 0.31 5.10e-17 0.00 0.00 0.28 0.00 0.25 0.33 0.00 0.26 0.43 0.00
+GD 0.20 0.00 0.52 4.78e-12 0.25 0.12 0.46 0.02 0.35 0.50 0.28 0.42 0.50 0.13
+IHL 0.41 0.38 0.51 1.46e-14 0.77 0.85 0.49 0.36 0.38 0.46 0.69 0.42 0.52 0.50
+GD+FILA 0.08 0.00 0.50 1.16e-05 0.09 0.00 0.22 0.00 0.31 0.52 0.04 0.26 0.34 0.00
+IHL+FILA 0.13 0.03 0.56 1.21e-01 0.70 0.80 0.47 0.32 0.40 0.48 0.67 0.44 0.55 0.50
+Table 4: Detailed TOFU Results using the Phi-1.5B model with LoRA rank = 16.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.93 0.93 0.48 1.15e-17 0.92 0.92 0.48 0.41 0.37 0.45 0.75 0.41 0.50 0.52
+Retain90 0.43 0.14 0.63 1.00e+00 0.91 0.92 0.48 0.43 0.38 0.45 0.76 0.41 0.49 0.52
+TOFU Forget01
+KL 0.68 0.72 0.48 4.17e-05 0.87 0.90 0.49 0.43 0.37 0.45 0.77 0.41 0.49 0.52
+DPO 0.84 0.90 0.51 4.72e-04 0.87 0.91 0.47 0.43 0.38 0.45 0.76 0.42 0.52 0.52
+NPO 0.65 0.71 0.49 5.05e-05 0.87 0.89 0.48 0.42 0.37 0.44 0.75 0.41 0.49 0.51
+GA 0.67 0.72 0.49 5.56e-05 0.87 0.89 0.48 0.42 0.37 0.45 0.75 0.41 0.49 0.51
+GA+FILA 0.03 0.00 0.78 5.55e-06 0.02 0.00 0.16 0.00 0.26 0.27 0.01 0.26 0.28 0.00
+GD 0.68 0.78 0.48 8.87e-05 0.90 0.92 0.49 0.40 0.38 0.45 0.75 0.41 0.50 0.52
+IHL 0.65 0.77 0.48 1.28e-04 0.90 0.92 0.49 0.42 0.38 0.45 0.76 0.41 0.50 0.52
+GD+FILA 0.04 0.00 0.77 1.15e-03 0.03 0.00 0.17 0.00 0.26 0.24 0.02 0.25 0.26 0.00
+IHL+FILA 0.37 0.09 0.61 3.06e-02 0.71 0.82 0.49 0.43 0.39 0.47 0.73 0.44 0.53 0.52
+TOFU Forget05
+KL 0.00 0.00 0.76 4.87e-12 0.01 0.00 0.16 0.00 0.26 0.26 0.00 0.27 0.26 0.00
+DPO 0.35 0.85 0.49 3.17e-15 0.76 0.89 0.47 0.34 0.37 0.43 0.72 0.41 0.50 0.49
+NPO 0.45 0.17 0.61 3.64e-05 0.46 0.19 0.38 0.37 0.33 0.40 0.68 0.36 0.43 0.36
+GA 0.00 0.00 0.76 2.17e-13 0.01 0.00 0.16 0.00 0.26 0.26 0.00 0.26 0.25 0.00
+GA+FILA 0.00 0.00 0.22 4.77e-17 0.00 0.00 0.35 0.00 0.22 0.35 0.00 0.19 0.37 0.00
+GD 0.24 0.03 0.56 1.76e-03 0.32 0.18 0.44 0.06 0.35 0.41 0.39 0.35 0.43 0.23
+IHL 0.45 0.42 0.50 4.18e-11 0.79 0.87 0.49 0.38 0.38 0.46 0.71 0.41 0.50 0.51
+GD+FILA 0.04 0.00 0.71 4.16e-06 0.05 0.00 0.17 0.00 0.26 0.23 0.02 0.27 0.28 0.00
+IHL+FILA 0.34 0.04 0.60 3.02e-03 0.71 0.81 0.48 0.37 0.38 0.46 0.69 0.42 0.52 0.50
+TOFU Forget10
+KL 0.01 0.00 0.60 2.17e-06 0.01 0.00 0.17 0.00 0.29 0.43 0.00 0.25 0.40 0.00
+DPO 0.28 0.85 0.48 2.51e-18 0.81 0.91 0.48 0.32 0.37 0.43 0.71 0.41 0.49 0.49
+NPO 0.44 0.25 0.65 2.31e-03 0.45 0.28 0.35 0.39 0.33 0.38 0.67 0.37 0.42 0.38
+GA 0.01 0.00 0.60 2.17e-06 0.01 0.00 0.17 0.00 0.28 0.42 0.00 0.25 0.39 0.00
+GA+FILA 0.00 0.00 0.23 4.22e-21 0.00 0.00 0.33 0.00 0.29 0.53 0.00 0.27 0.44 0.00
+GD 0.11 0.00 0.45 3.33e-06 0.39 0.34 0.42 0.09 0.36 0.53 0.34 0.41 0.53 0.29
+IHL 0.34 0.20 0.53 2.89e-11 0.81 0.86 0.50 0.42 0.39 0.47 0.70 0.42 0.53 0.52
+GD+FILA 0.10 0.00 0.43 2.02e-08 0.10 0.00 0.27 0.00 0.25 0.38 0.03 0.27 0.40 0.00
+IHL+FILA 0.13 0.02 0.68 2.08e-02 0.66 0.79 0.46 0.42 0.38 0.46 0.72 0.44 0.52 0.51
+Table 5: Detailed TOFU Results using the Phi-1.5B model with LoRA rank = 32.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.99 0.99 0.51 2.19e-20 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.55 0.63
+Retain90 0.40 0.15 0.67 1.00e+00 0.98 0.99 0.47 0.92 0.47 0.61 0.88 0.42 0.55 0.63
+TOFU Forget01
+KL 0.95 0.99 0.55 9.73e-05 0.98 0.99 0.47 0.94 0.47 0.62 0.90 0.43 0.56 0.63
+DPO 0.95 0.99 0.56 1.40e-04 0.98 0.99 0.47 0.93 0.47 0.62 0.89 0.43 0.55 0.63
+NPO 0.95 0.99 0.55 9.73e-05 0.98 0.99 0.47 0.93 0.47 0.62 0.89 0.43 0.56 0.63
+GA 0.95 0.99 0.55 6.71e-05 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.56 0.63
+GA+FILA 0.03 0.00 0.87 3.12e-05 0.04 0.00 0.12 0.01 0.25 0.22 0.01 0.27 0.25 0.00
+GD 0.95 0.99 0.55 1.40e-04 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.55 0.63
+IHL 0.95 0.99 0.55 1.17e-04 0.98 0.99 0.47 0.94 0.47 0.62 0.90 0.43 0.55 0.63
+GD+FILA 0.03 0.00 0.87 1.15e-05 0.04 0.00 0.13 0.00 0.25 0.21 0.02 0.27 0.24 0.00
+IHL+FILA 0.69 0.84 0.55 1.53e-04 0.98 0.99 0.47 0.93 0.46 0.60 0.89 0.42 0.54 0.62
+TOFU Forget05
+KL 0.92 0.96 0.53 9.25e-17 0.97 0.98 0.46 0.93 0.48 0.63 0.90 0.45 0.57 0.64
+DPO 0.83 0.95 0.57 8.99e-14 0.86 0.95 0.44 0.92 0.47 0.60 0.87 0.45 0.56 0.62
+NPO 0.89 0.95 0.54 2.47e-16 0.95 0.97 0.46 0.94 0.48 0.63 0.90 0.45 0.57 0.64
+GA 0.90 0.96 0.54 6.50e-16 0.96 0.98 0.46 0.94 0.48 0.63 0.90 0.45 0.57 0.64
+GA+FILA 0.01 0.00 0.83 1.23e-15 0.01 0.00 0.10 0.00 0.24 0.17 0.00 0.26 0.24 0.00
+GD 0.93 0.97 0.52 6.50e-16 0.98 0.99 0.47 0.94 0.48 0.62 0.89 0.44 0.56 0.64
+IHL 0.94 0.97 0.52 6.64e-17 0.98 0.99 0.47 0.94 0.47 0.62 0.90 0.44 0.56 0.64
+GD+FILA 0.02 0.00 0.77 1.50e-08 0.03 0.00 0.14 0.01 0.24 0.17 0.00 0.24 0.21 0.00
+IHL+FILA 0.54 0.70 0.58 6.87e-13 0.90 0.94 0.45 0.92 0.48 0.62 0.89 0.47 0.60 0.64
+TOFU Forget10
+KL 0.47 0.32 0.65 2.56e-05 0.47 0.33 0.35 0.93 0.43 0.55 0.89 0.43 0.56 0.49
+DPO 0.45 0.94 0.55 5.10e-17 0.66 0.96 0.44 0.82 0.43 0.54 0.87 0.42 0.51 0.57
+NPO 0.54 0.31 0.65 3.33e-06 0.54 0.31 0.35 0.94 0.39 0.50 0.90 0.39 0.51 0.47
+GA 0.49 0.15 0.66 2.31e-03 0.49 0.15 0.33 0.93 0.39 0.51 0.91 0.37 0.50 0.39
+GA+FILA 0.02 0.00 0.86 5.40e-18 0.02 0.00 0.09 0.00 0.24 0.19 0.00 0.24 0.18 0.00
+GD 0.82 0.92 0.51 2.19e-16 0.92 0.97 0.47 0.92 0.46 0.60 0.88 0.42 0.55 0.62
+IHL 0.73 0.87 0.57 3.71e-15 0.88 0.93 0.45 0.94 0.49 0.64 0.89 0.47 0.59 0.64
+GD+FILA 0.01 0.00 0.85 1.83e-21 0.01 0.00 0.09 0.00 0.25 0.18 0.00 0.23 0.18 0.00
+IHL+FILA 0.30 0.28 0.65 2.93e-01 0.91 0.96 0.45 0.89 0.47 0.62 0.88 0.45 0.57 0.63
+Table 6: Detailed TOFU Results using the Llama2-7B model with LoRA rank = 4.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.99 0.99 0.51 2.19e-20 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.55 0.63
+Retain90 0.40 0.15 0.67 1.00e+00 0.98 0.99 0.47 0.92 0.47 0.61 0.88 0.42 0.55 0.63
+TOFU Forget01
+KL 0.95 0.98 0.56 1.17e-04 0.98 0.99 0.47 0.93 0.47 0.62 0.90 0.43 0.56 0.64
+DPO 0.94 0.99 0.56 1.40e-04 0.98 0.99 0.47 0.94 0.47 0.62 0.90 0.43 0.56 0.63
+NPO 0.94 0.98 0.55 6.71e-05 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.56 0.63
+GA 0.94 0.98 0.55 6.71e-05 0.98 0.99 0.47 0.93 0.47 0.62 0.89 0.43 0.55 0.63
+GA+FILA 0.01 0.00 0.86 3.12e-05 0.02 0.00 0.09 0.01 0.25 0.22 0.00 0.24 0.17 0.00
+GD 0.95 0.99 0.55 1.17e-04 0.98 0.99 0.47 0.93 0.47 0.62 0.89 0.43 0.56 0.63
+IHL 0.95 0.99 0.55 1.40e-04 0.98 0.99 0.47 0.94 0.47 0.61 0.90 0.43 0.55 0.63
+GD+FILA 0.02 0.00 0.85 2.35e-06 0.02 0.00 0.09 0.00 0.25 0.22 0.00 0.24 0.19 0.00
+IHL+FILA 0.57 0.68 0.55 2.60e-04 0.97 0.98 0.47 0.92 0.46 0.60 0.89 0.42 0.54 0.62
+TOFU Forget05
+KL 0.77 0.89 0.54 6.50e-16 0.88 0.95 0.45 0.94 0.47 0.61 0.91 0.45 0.57 0.63
+DPO 0.14 0.90 0.58 6.87e-13 0.27 0.91 0.43 0.59 0.44 0.57 0.84 0.44 0.54 0.50
+NPO 0.62 0.72 0.57 1.09e-14 0.68 0.76 0.44 0.87 0.46 0.61 0.88 0.46 0.57 0.60
+GA 0.64 0.78 0.56 1.21e-13 0.71 0.82 0.44 0.89 0.47 0.62 0.90 0.46 0.57 0.61
+GA+FILA 0.01 0.00 0.88 1.32e-19 0.01 0.00 0.07 0.01 0.25 0.19 0.00 0.24 0.17 0.00
+GD 0.90 0.94 0.52 3.17e-15 0.97 0.98 0.47 0.92 0.48 0.62 0.89 0.43 0.56 0.64
+IHL 0.84 0.94 0.53 1.24e-17 0.98 0.98 0.47 0.94 0.47 0.61 0.90 0.44 0.56 0.63
+GD+FILA 0.02 0.00 0.86 4.46e-18 0.03 0.00 0.09 0.01 0.25 0.19 0.01 0.24 0.19 0.00
+IHL+FILA 0.40 0.39 0.62 4.77e-07 0.86 0.94 0.45 0.92 0.48 0.63 0.91 0.45 0.59 0.63
+TOFU Forget10
+KL 0.31 0.00 0.65 7.31e-03 0.31 0.00 0.22 0.07 0.29 0.40 0.62 0.34 0.48 0.01
+DPO 0.34 0.94 0.54 5.10e-17 0.67 0.96 0.45 0.65 0.42 0.53 0.84 0.40 0.49 0.55
+NPO 0.50 0.26 0.67 1.73e-05 0.51 0.26 0.33 0.91 0.39 0.51 0.91 0.38 0.49 0.45
+GA 0.11 0.00 0.65 3.32e-08 0.12 0.00 0.17 0.00 0.27 0.38 0.10 0.28 0.42 0.00
+GA+FILA 0.00 0.00 0.61 3.32e-08 0.00 0.00 0.15 0.00 0.26 0.36 0.00 0.21 0.24 0.00
+GD 0.47 0.47 0.49 3.71e-15 0.52 0.67 0.48 0.87 0.44 0.58 0.87 0.40 0.56 0.56
+IHL 0.59 0.77 0.58 2.86e-14 0.86 0.91 0.45 0.93 0.51 0.66 0.89 0.48 0.62 0.65
+GD+FILA 0.03 0.00 0.85 1.15e-17 0.04 0.00 0.07 0.01 0.25 0.21 0.01 0.23 0.16 0.00
+IHL+FILA 0.15 0.11 0.74 3.63e-07 0.93 0.97 0.44 0.86 0.49 0.64 0.86 0.45 0.59 0.63
+Table 7: Detailed TOFU Results using the Llama2-7B model with LoRA rank = 8.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.99 0.99 0.51 2.19e-20 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.55 0.63
+Retain90 0.40 0.15 0.67 1.00e+00 0.98 0.99 0.47 0.92 0.47 0.61 0.88 0.42 0.55 0.63
+TOFU Forget01
+KL 0.90 0.96 0.56 1.07e-04 0.98 0.99 0.47 0.93 0.48 0.62 0.89 0.44 0.56 0.64
+DPO 0.94 0.99 0.56 3.37e-04 0.97 0.99 0.46 0.94 0.48 0.62 0.90 0.44 0.55 0.63
+NPO 0.89 0.96 0.55 8.09e-05 0.98 0.99 0.47 0.93 0.47 0.62 0.89 0.43 0.56 0.64
+GA 0.90 0.96 0.55 9.73e-05 0.98 0.99 0.47 0.93 0.48 0.62 0.90 0.43 0.56 0.64
+GA+FILA 0.00 0.00 0.69 1.58e-02 0.01 0.00 0.17 0.00 0.27 0.24 0.01 0.26 0.21 0.00
+GD 0.92 0.97 0.55 1.28e-04 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.56 0.63
+IHL 0.88 0.97 0.55 1.40e-04 0.98 0.99 0.47 0.93 0.47 0.61 0.90 0.43 0.55 0.63
+GD+FILA 0.02 0.00 0.67 3.33e-01 0.01 0.00 0.18 0.00 0.27 0.27 0.01 0.28 0.25 0.00
+IHL+FILA 0.52 0.42 0.56 7.13e-04 0.91 0.97 0.47 0.89 0.46 0.61 0.88 0.42 0.55 0.62
+TOFU Forget05
+KL 0.51 0.52 0.58 2.80e-12 0.54 0.58 0.43 0.90 0.41 0.52 0.91 0.42 0.53 0.54
+DPO 0.17 0.91 0.57 1.62e-13 0.44 0.93 0.44 0.60 0.43 0.55 0.84 0.42 0.51 0.53
+NPO 0.47 0.48 0.63 8.80e-10 0.48 0.52 0.39 0.89 0.44 0.57 0.90 0.45 0.57 0.53
+GA 0.48 0.45 0.62 1.12e-09 0.50 0.50 0.39 0.90 0.41 0.52 0.90 0.43 0.54 0.52
+GA+FILA 0.03 0.00 0.68 1.00e-03 0.03 0.00 0.15 0.01 0.24 0.22 0.02 0.22 0.19 0.00
+GD 0.68 0.79 0.51 4.96e-14 0.82 0.94 0.47 0.90 0.46 0.59 0.86 0.41 0.56 0.61
+IHL 0.69 0.84 0.54 3.42e-16 0.94 0.97 0.46 0.94 0.46 0.59 0.90 0.44 0.56 0.63
+GD+FILA 0.05 0.00 0.80 4.19e-10 0.06 0.00 0.12 0.01 0.26 0.24 0.02 0.26 0.22 0.00
+IHL+FILA 0.25 0.18 0.69 1.34e-02 0.88 0.95 0.46 0.93 0.47 0.62 0.89 0.44 0.58 0.63
+TOFU Forget10
+KL 0.00 0.00 0.81 2.86e-14 0.00 0.00 0.09 0.00 0.22 0.23 0.00 0.23 0.22 0.00
+DPO 0.27 0.94 0.53 1.06e-16 0.73 0.97 0.45 0.58 0.40 0.52 0.85 0.39 0.48 0.54
+NPO 0.47 0.26 0.70 1.66e-04 0.47 0.27 0.31 0.91 0.41 0.53 0.90 0.38 0.50 0.44
+GA 0.00 0.00 0.79 1.60e-11 0.00 0.00 0.11 0.00 0.22 0.23 0.00 0.22 0.24 0.00
+GA+FILA 0.00 0.00 0.90 4.07e-32 0.00 0.00 0.04 0.00 0.25 0.17 0.00 0.24 0.13 0.00
+GD 0.39 0.06 0.47 7.41e-13 0.44 0.19 0.52 0.81 0.42 0.58 0.86 0.38 0.55 0.44
+IHL 0.43 0.52 0.63 5.73e-07 0.89 0.94 0.45 0.90 0.52 0.67 0.89 0.48 0.62 0.65
+GD+FILA 0.04 0.00 0.70 2.65e-02 0.04 0.00 0.16 0.00 0.27 0.31 0.02 0.23 0.27 0.00
+IHL+FILA 0.04 0.02 0.79 8.69e-10 0.91 0.97 0.44 0.88 0.50 0.65 0.87 0.46 0.59 0.64
+Table 8: Detailed TOFU Results using the Llama2-7B model with LoRA rank = 16.
+Method Forget Quality Model Utility
+Forget Set FQ ($\uparrow$) Retain Set Real Authors Real World MU ($\uparrow$)
+Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio Rouge-L Prob. Truth Ratio
+Original 0.99 0.99 0.51 2.19e-20 0.98 0.99 0.47 0.94 0.47 0.62 0.89 0.43 0.55 0.63
+Retain90 0.40 0.15 0.67 1.00e+00 0.98 0.99 0.47 0.92 0.47 0.61 0.88 0.42 0.55 0.63
+TOFU Forget01
+KL 0.78 0.90 0.56 5.05e-05 0.96 0.98 0.47 0.93 0.48 0.62 0.90 0.44 0.56 0.64
+DPO 0.88 0.97 0.58 6.57e-04 0.90 0.97 0.46 0.93 0.48 0.63 0.90 0.45 0.56 0.63
+NPO 0.73 0.90 0.55 1.40e-04 0.95 0.98 0.47 0.93 0.48 0.63 0.90 0.44 0.56 0.64
+GA 0.76 0.90 0.56 8.87e-05 0.96 0.98 0.47 0.93 0.47 0.62 0.90 0.44 0.56 0.63
+GA+FILA 0.03 0.00 0.77 7.70e-03 0.02 0.00 0.13 0.00 0.25 0.21 0.00 0.23 0.18 0.00
+GD 0.82 0.91 0.55 1.53e-04 0.96 0.98 0.47 0.93 0.47 0.62 0.88 0.43 0.56 0.63
+IHL 0.76 0.89 0.55 2.18e-04 0.96 0.98 0.47 0.93 0.47 0.61 0.90 0.43 0.55 0.63
+GD+FILA 0.04 0.00 0.81 1.67e-04 0.03 0.00 0.14 0.00 0.24 0.21 0.01 0.25 0.19 0.00
+IHL+FILA 0.43 0.20 0.61 4.13e-03 0.81 0.94 0.47 0.89 0.45 0.60 0.88 0.42 0.55 0.61
+TOFU Forget05
+KL 0.45 0.09 0.64 4.81e-04 0.46 0.12 0.36 0.91 0.34 0.45 0.90 0.36 0.50 0.36
+DPO 0.30 0.93 0.56 2.72e-14 0.77 0.96 0.44 0.78 0.42 0.54 0.86 0.41 0.50 0.57
+NPO 0.46 0.29 0.70 1.81e-05 0.47 0.33 0.32 0.94 0.40 0.51 0.91 0.43 0.54 0.47
+GA 0.44 0.09 0.69 1.14e-01 0.44 0.11 0.29 0.85 0.37 0.49 0.92 0.37 0.50 0.34
+GA+FILA 0.01 0.00 0.86 2.47e-16 0.02 0.00 0.09 0.00 0.24 0.20 0.01 0.24 0.17 0.00
+GD 0.40 0.30 0.50 6.87e-13 0.49 0.62 0.47 0.82 0.41 0.53 0.86 0.41 0.56 0.54
+IHL 0.51 0.63 0.56 2.12e-12 0.88 0.94 0.46 0.93 0.47 0.60 0.90 0.44 0.57 0.63
+GD+FILA 0.11 0.00 0.68 1.00e-03 0.11 0.00 0.18 0.01 0.24 0.27 0.03 0.21 0.23 0.00
+IHL+FILA 0.20 0.05 0.75 7.38e-03 0.87 0.95 0.48 0.91 0.45 0.59 0.88 0.44 0.58 0.62
+TOFU Forget10
+KL 0.00 0.00 0.86 4.22e-21 0.00 0.00 0.07 0.00 0.26 0.25 0.00 0.25 0.23 0.00
+DPO 0.23 0.94 0.52 5.10e-17 0.88 0.98 0.46 0.75 0.40 0.50 0.86 0.38 0.47 0.56
+NPO 0.47 0.28 0.71 4.69e-04 0.48 0.31 0.31 0.87 0.43 0.56 0.89 0.40 0.52 0.46
+GA 0.00 0.00 0.60 5.07e-06 0.00 0.00 0.20 0.00 0.19 0.28 0.00 0.20 0.35 0.00
+GA+FILA 0.00 0.00 0.67 3.33e-06 0.00 0.00 0.18 0.00 0.26 0.32 0.00 0.31 0.39 0.00
+GD 0.01 0.00 0.56 5.44e-08 0.18 0.20 0.65 0.43 0.48 0.63 0.81 0.43 0.58 0.38
+IHL 0.14 0.11 0.79 1.22e-08 0.94 0.97 0.45 0.88 0.48 0.62 0.88 0.45 0.58 0.63
+GD+FILA 0.03 0.00 0.83 9.16e-16 0.04 0.00 0.11 0.00 0.24 0.20 0.01 0.25 0.21 0.00
+IHL+FILA 0.04 0.02 0.73 3.77e-05 0.90 0.96 0.48 0.92 0.52 0.67 0.87 0.46 0.60 0.65
+Table 9: Detailed TOFU Results using the Llama2-7B model with LoRA rank = 32.

### v5

@@
 We find that combining GA with low-rank adaptation results in poor trade-offs between computational cost and generative performance.
-To address these challenges, we propose two novel techniques for robust and efficient unlearning for LLMs.
+To address these challenges, we propose Low-rank Knowledge Unlearning (LoKU), a novel framework that enables robust and efficient unlearning for LLMs.
 First, we introduce Inverted Hinge Loss, which suppresses unwanted tokens while maintaining fluency by boosting the probability of the next most likely token.
@@
 However, the empirical effects of LoRA in the context of LLM unlearning remain largely unexplored.
-In this paper, we present the first in-depth study of LLM unlearning under the low-rank adaptation paradigm and introduce two novel techniques for robust and parameter-efficient knowledge unlearning.
+In this paper, we present the first in-depth study of LLM unlearning under the low-rank adaptation paradigm and introduce Low-rank Knowledge Unlearning (LoKU), which consists of two novel techniques for robust and parameter-efficient knowledge unlearning.
 First, we analyze the derivatives of GA to highlight its drawbacks such as (1) unnecessary forgetting induced by indiscriminately increasing the probability of all tokens and (2) unstable optimization due to the unboundedness of maximizing the next-token prediction loss.
@@
   • We introduce FILA, a method to accelerate unlearning by data-adaptively assigning parameters responsible for unwanted information to low-rank adapters before tuning.
-  • We demonstrate that IHL combined with FILA outperforms previous baselines in terms of both efficiency and post-unlearning performance (Figure 2).
+  • We demonstrate that our proposed LoKU using both IHL and FILA outperforms previous baselines in terms of both efficiency and post-unlearning performance (Figure 2).
 Figure 2: Compute cost for successful unlearning vs.
@@
 Unlearning via gradient differences (GD) with vanilla LoRA leads to significant loss in performance compared to full-parameter GD unlearning due to lack of plasticity.
-However, our proposed method using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the parameter-efficiency of LoRA.
+However, our proposed LoKU using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the parameter-efficiency of LoRA.
 2 Related Work
@@
 These areas have not been adequately addressed in previous studies, highlighting the contributions of our work.
-3 Proposed Method
+3 Proposed Method: Low-rank Knowledge Unlearning (LoKU)
 3.1 Preliminaries
@@
 Intuitively, our Fisher-weighted Initialization of Low-rank Adapters (FILA) extracts parameters that are important for generating ${\mathcal{D}}_{f}$ but not for generating ${\mathcal{D}}_{r}$, such that LoRA tuning can be focused on erasing knowledge relevant to ${\mathcal{D}}_{f}$ while keeping information regarding ${\mathcal{D}}_{r}$.
-3.5 Final Loss Function for LLM Unlearning
+3.5 Final Loss Function of LoKU
 In summary, we perform unlearning on the model $\bm{\Theta}=\bm{\theta}\cup\bm{\theta}_{\text{FILA}}$ consisted of original pretrained weights $\bm{\theta}$ and FILA-initialized low-rank adapter weights for each linear layer $\bm{\theta}_{\text{FILA}}=\{{\bm{A}}^{*}_{\ell},{\bm{B}}^{*}_{\ell}\}_{\ell=1}^{L}$, where $L$ represents the number of layers tuned via LoRA.
 Additionally, we incorporate GD, which utilizes the auxiliary retain set $\mathcal{D}_{r}$.
-The final loss function using both the proposed IHL and FILA is defined as follows:
+The final loss function of LoKU is defined as follows:
 $\operatorname*{minimize}_{\theta_{\text{FILA}}}\sum_{{\bm{x}}_{r}\in{\mathcal{D}}_{f},{\bm{x}}_{f}\in{\mathcal{D}}_{r}}\mathcal{L}_{\text{IHL}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (3)
@@
 GD+FILA 7.4 1.2 27.4 42.0 (-1.4) 6.5 (-2.9) 89.5 (+71.7)
-IHL+FILA 6.0 0.3 23.9 42.2 (-1.2) 10.1 (+0.7) 24.0 (+6.2)
+LoKU (IHL+FILA) 6.0 0.3 23.9 42.2 (-1.2) 10.1 (+0.7) 24.0 (+6.2)
 GPT-Neo 1.3B Before - - 67.6 92.2 49.8 11.5 11.5
@@
 GD+FILA 7.8 1.9 23.2 44.2 (-5.6) 5.5 (-6.0) 54.5 (+43.0)
-IHL+FILA 13.0 0.5 29.6 48.3 (-1.5) 12.1 (+0.6) 14.7 (+3.2)
+LoKU (IHL+FILA) 13.0 0.5 29.6 48.3 (-1.5) 12.1 (+0.6) 14.7 (+3.2)
 GPT-Neo 2.7B Before - - 70.4 93.4 52.3 11.5 10.4
@@
 GD+FILA 6.8 1.6 28.9 44.8 (-7.5) 9.3 (-2.2) 68.7 (+58.3)
-IHL+FILA 10.3 0.1 28.5 49.6 (-2.7) 10.7 (-0.8) 16.0 (+5.6)
+LoKU (IHL+FILA) 10.3 0.1 28.5 49.6 (-2.7) 10.7 (-0.8) 16.0 (+5.6)
 4.1 Training Data Extraction Challenge
@@
 Running GD with FILA notably increases the rate of unlearning success, but at significant cost in overall performance.
-Using both IHL and FILA best minimizes post-unlearning performance degradation in all three aspects.
+Our LoKU using both IHL and FILA best minimizes post-unlearning performance degradation in all aspects.
 Results.
@@
 In essence, whether the ability of FILA to accelerate tuning also translates to benefits in downstream performance depends on the loss function being optimized.
-On the other hand, when paired with our bounded IHL, FILA leads to better retention of LLM performance, showcasing the strong synergy of our approach that enjoys not only the stability of IHL but also the speedup from FILA.
+On the other hand, when paired with our bounded IHL (i.e., LoKU), FILA leads to better retention of LLM performance, showcasing the strong synergy of our approach that enjoys not only the stability of IHL but also the speedup from FILA.
 4.2 Analysis
@@
 Notably, when LoRA is applied to FFNs, we observe significant increase in rate of successful unlearning.
-Furthermore, integrating FILA with IHL achieves the best post-unlearning performance across all LoRA target module combinations.
+Furthermore, our LoKU integrating FILA with IHL, achieves the best post-unlearning performance across all LoRA target module combinations.
 Figure 4: Results from unlearning examples from TDEC dataset using LoRA with rank 32 to adapt sets of layers on GPT-Neo-125M.
@@
 Both GD and GD+FILA suffers from significant loss of model utility due to using GA for unlearning.
-Replacing GA with IHL largely retains model utility, then initializing LoRA adapters with FILA significantly boosts the unlearning efficiency of IHL.
+Replacing GA with IHL largely retains model utility, then our LoKU initializing LoRA adapters with FILA significantly boosts the unlearning efficiency of IHL.
 Results.
@@
 This stability comes at the cost of unlearning efficiency, however, as randomly initialized LoRA weights are unable to effectively represent weight changes required to decrease IHL.
-Nonetheless, initializing LoRA adapters with FILA largely alleviates this issue and significantly enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.
+Nonetheless, our LoKU initializing LoRA adapters with FILA largely alleviates this issue and significantly enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.
 Interestingly, we find the prior weight assignment via FILA can lead to excessive unlearning in some cases (e.g., unlearning 10% forget set with ranks 8 or 16 on Llama2-7B), with model updates reducing the forget quality after reaching the upper bound at zero.
@@
 We also propose Fisher-weighted initialization for low-rank adaptation (FILA) that pre-assigns weights relatively important to generating unwanted information as means to facilitate efficient LLM unlearning with LoRA.
-Experiments on the Training Data Extraction Challenge dataset with GPT-Neo models along with the TOFU benchmark using Phi-1.5B and Llama2-7B models show that our proposed methods enable faster and more stable LoRA-based LLM unlearning, significantly outperforming existing baselines in computational efficiency as well as post-unlearning performance.
+Experiments on the Training Data Extraction Challenge dataset with GPT-Neo models along with the TOFU benchmark using Phi-1.5B and Llama2-7B models show that our proposed Low-rank Knowledge Unlearning (LoKU) enables faster and more stable LoRA-based LLM unlearning, significantly outperforming existing baselines in computational efficiency as well as post-unlearning performance.
 Appendix A Derivative Analysis for the Inverted Hinge Loss Function